### v3

# Tug-of-war between idiom’s figurative and literal meanings in LLMs

Abstract
Idioms present a unique challenge for language models due to their non-compositional figurative meanings, which often strongly diverge from the idiom’s literal interpretation. This duality requires a model to learn representing and deciding between the two meanings to interpret an idiom in a figurative sense, or literally. In this paper, we employ tools from mechanistic interpretability to trace how a large pretrained causal transformer (LLama3.2-1B-base) deals with this ambiguity. We localize three steps of idiom processing: First, the idiom’s figurative meaning is retrieved in early attention and MLP sublayers. We identify specific attention heads which boost the figurative meaning of the idiom while suppressing the idiom’s literal interpretation. The model subsequently represents the figurative representation through an intermediate path. Meanwhile, a parallel bypass route forwards literal interpretation, ensuring that a both reading remain available. Overall, our findings provide a mechanistic evidence for idiom comprehension in an autoregressive transformer. [^1] Code and dataset are available at https://github.com/sori424/idiom_processing under the MIT License.
$\dagger$\$$\dagger$\$footnotetext: Co-corresponding authors.
1 Introduction
Figure 1: The figurative and literal interpretations are highlighted in the block and path. We find three main steps for idiom processing: 1 Idiom retrieval step: Early layers (i.e., 0-3) attention and MLP are actively retrieving the idiom’s figurative meaning while storing both figurative and literal interpretations in the residual stream. 2 Selective interpretation step: At the token immediately following the idiom span, the model begins to encode a representation that favors the figurative interpretation over the literal one, starting from the middle layers. 3 Interpretation routing: For final prediction, the model passes literal interpretation via both a direct compositional semantic path ( 3-2), as well as the intermediate pathway that prioritizes the figurative meaning ( 3-1 figurative path).
Idioms challenge standard semantic composition because they are, by definition, multi-word sequences whose figurative meanings diverge from the composition of their literal parts Beck and Weber (2020). For instance, “kick the bucket” conveys the figurative meaning “to die”, yet its literal sense “physically kicking a container” remains semantically available. This ambiguity often leads large language models (LLMs) to misinterpret idioms, favoring their literal sense over the intended figurative one Kabra et al. (2023); Liu et al. (2023); Knietaite et al. (2024). Psycholinguistics has been investigated how humans understand idioms, whether they access the idiomatic meaning directly or first have to access the literal meaning of the idiom and only in a later stage suppress that meaning and access the figurative meaning. These findings have lead to various models describing human idiom processing Gibbs (1980); Cacciari and Glucksberg (1991); Bobrow and Bell (1973); Cacciari and Tabossi (1988).
Yet for LLMs, it’s unclear how they process idioms, potentially limiting further improvements in their idiom comprehension. One explanation may lie in how the model transforms raw input token embeddings into richer representations Tian et al. (2023); Haviv et al. (2022); Dankers et al. (2022). Within the autoregressive transformer, token representations are incrementally refined from the initial input embeddings to the final representation used for next token prediction Vig et al. (2020). This transformation is achieved through a series of residual layers, which consist of multi-head self-attention (MHSA) and MLP sublayers, each adding their outputs to update the representation. This multi-stage transformation suggests that idiom processing might follow a multi-step transformation process. Haviv et al. (2022) demonstrated that the early layers of transformer models store and retrieve idiomatic information, while the higher layers enhance model confidence. Meanwhile, Dankers et al. (2022) showed that during machine translation of idioms, the transformer model recognizes idiomatic expressions and processes them more as single lexical units. This involves grouping the idioms together using attention sublayers and reducing their interaction with surrounding context. However, both studies lack a detailed mapping of the specific model components and causal pathways underlying these idiom processing behaviors.
In this paper, we first aim to identify how the idiom is interpreted by the transformer (i.e., LLama3.2-1B-base), where its figurative meaning is located, and how it is retrieved by the model. To this end, we pinpoint specific components of the model that have specialized in processing idioms by boosting their figurative meaning, while suppressing the compositional semantic literal interpretation. Having identified these, we proceed to investigating the information flow in the model, with the goal of identifying specific pathways through which the figurative meaning is passed forward to the final prediction, and whether this differs from the pathways of passing forward information about an idiom’s literal interpretation.
We address these questions through knockout analyses Nanda et al. (2023); Wang et al. (2022); Geva et al. (2023), where we separately knock out activations of each component (layer-wise MHSA and MLP, individual attention heads) in the model to observe their importance for retrieving idiom’s figurative meaning. Furthermore, to trace the flow of these interpretations, we employ activation patching experiments Wang et al. (2022); Meng et al. (2022); Hanna et al. (2023); Conmy et al. (2023); Stolfo et al. (2023), in which we ablate one or both interpretations from specific pathways within the model.
Our main findings are summarized in Figure 1: we identify three main steps of idiom processing (1) retrieval of the figurative meaning of the idiom in the early layers, (2) selective encoding of the two interpretations at intermediate layers after the idiom span token position, and (3) two alternative paths for routing the idiom’s meaning to the final prediction, via a direct path for the literal meaning and an intermediate pathway that preferentially amplifies figurative interpretation over the other. In summary, our main contributions are:
  • We propose a next token prediction task that distinguishes figurative vs. literal interpretations using interpretation-specific tokens (Section 2).
  • We show that early layers contribute to retrieve the idiom’s figurative meaning, and that these are significantly different from general semantic processing (Section 4).
  • We identify a selection mechanism at the token immediately following the idiom, where its residual activation shifts toward the figurative interpretation and then propagates through the intermediate layers (Section 5).
  • We also uncover a bypass pathway that channels the literal interpretation directly into the final prediction, circumventing the intermediate route (Section 5.2).
2 Idiom interpretation task
We are interested in teasing apart the language model’s processing of the literal vs. figurative interpretation of ambiguous idiom expressions. We therefore formalize a task and construct a dataset to investigate the interpretation shift in the model. In order to allow for effective probing, we embed the idiom into a sentence template, constructed such that the interpretation is revealed via next token prediction. The template [^2] X is instantiated with various pronouns (he/she/it/they); the word “would” is inserted only in some of the idioms to make the sentence sound more fluent; idioms are morphologically fit into the sentence; depending on sg/pl form of X, the second part of the template contained was or were. ‘X (would) IDIOM because X was so/too/a/the’ can be applied to all of the selected idioms while preserving the idiom’s ambiguity. It elicits a causal completion, which we expect to differ between the literal and the figurative meaning of the idiom: for the figurative meaning of the idiom “kick the bucket”, we would expect completions that provide causes for the figurative meaning “to die”, e.g., old/sick, while for the literal meaning, we would expect completions that provide reasons for “kicking an object”, e.g. angry/mad. Figure 2 illustrates this process and selected examples can be found in Appendix Table 2.
Figure 2: Pipeline for extracting figurative ($C_{f}$) and literal ($C_{l}$) candidate tokens for an idiomatic sentence.
Since there is no single ground truth continuation that reveals either the figurative or the literal reading, we automatically construct sets of multiple tokens that are consistent with either the literal or the figurative meaning. The tokens in each set are mutually exclusive and selected based on their high likelihood of being associated with either the figurative or literal meaning of the idiom. In the following, we describe each step of the process of generating the dataset.
1. Idiom Extraction: We select 245 idioms from a psycholinguistic paper Cronk et al. (1993), each of which received high literality ratings in a human rating study.
2. Sentence Generation: We generate a literal paraphrase and a figurative paraphrase for each idiom using the LLama3.3-70B-instruct model Grattafiori et al. (2024). The sentences below are three examples of the generated ambiguous sentence ($s_{a}$), figurative paraphrase ($s_{f}$), and literal paraphrase ($s_{l}$).
[…]
.[$s_{a}$] He would kick the bucket because he was so
[…]
.[$s_{f}$] He would die because he was so
[…]
.[$s_{l}$] He would kick the container because he was so
3. Token Set Generation: To generate sets of next-word completions that are indicative of the figurative vs. the literal interpretation, we prompt the LLaMa3.3-70B model with disambiguated paraphrases ($s_{f}$, $s_{l}$) and obtain logits from the model with greedy decoding. Let $z^{(f)}$ and $z^{(l)}$ denote the logits corresponding to $s_{f}$ and $s_{l}$ as input, respectively. We then compute the element-wise difference between the two sets of logits as follows:
$\Delta z_{v}=z_{v}^{(f)}-z_{v}^{(l)},$
where $v\in V$ is the vocabulary size of the model (i.e., $V=$128,256). Since tokens that score high in one context tend to score low in the other, $\Delta z_{v}$ captures each token’s context-specific relevance. The 20 tokens that have highest $\Delta z_{k}$ are chosen as the set $C_{f}$ of tokens that indicate a figurative interpretation and the 20 tokens that have lowest $\Delta z_{k}$ values are chosen as the set $C_{l}$ that indicates a literal interpretation.
4. Validation: To make sure that the Llama3.2-1B-base model has prior knowledge of the selected idioms, and the capability of discerning the context of the given sentence, we filter out the generated instances that satisfy the following two conditions:
(1) Given $s_{a}$ as a prompt, the cumulative probabilities of predicting a figurative candidate as a next token must exceed that of the literal candidate. This metric quantifies that the model has the knowledge of the idiom:
$\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{a})>\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{a}).$
(2) Given $s_{f}$ and $s_{l}$ as prompt, the cumulative probabilities of predicting next token candidate matching the context exceeds that of the opposite context:
$\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{f})>\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{f}),$
$\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{l})>\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{l}).$
110 out of the original set of 245 idioms passed these criteria. All following analyses are based on this subset of 110 validated idioms.
3 Methodology
3.1 Information flow in transformers
Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{T}]$ over a vocabulary $V$, each token is embedded as $\mathrm{x}_{i}^{0}=\mathrm{RoPE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{RoPE}$ applies rotary positional encoding to inject position $i$ Su et al. (2024), and $d$ is the model’s hidden dimension. From layer 0, these vectors $\mathrm{x}_{i}^{0}$ are carried forward and accumulated over the subsequent $L$ layers. At each layer $\ell$, the hidden state for token $i$ is updated from its previous value $\mathrm{x}_{i}^{\ell-1}$ by adding the multi-head self-attention (MHSA) output $a_{i}^{\ell}$ and MLP output $m_{i}^{\ell}$:
$\mathrm{x}_{i}^{\ell}=\mathrm{x}_{i}^{\ell-1}+a_{i}^{\ell}+m_{i}^{\ell}.$
MLP SublayersEvery MLP sublayer computes a local update for each representation:
$m_{i}^{\ell}=W_{F}^{\ell}\,\sigma\Bigl{(}W_{I}^{\ell}(a_{i}^{\ell}+\mathrm{x}_{i}^{\ell-1})\Bigr{)},$
where $W_{I}^{\ell}\in\mathbb{R}^{d_{ff}\times d}$ and $W_{F}^{\ell}\in\mathbb{R}^{d\times d_{ff}}$ are parameter matrices with inner-dimension $d_{ff}$ and $\sigma$ is a nonlinear activation function.
MHSA SublayersEach attention layer’s output $\mathrm{a}^{\ell}_{i}$ aggregates information of multiple parallel attention heads $h$. The parameter matrices of each MHSA sublayer comprise three projection matrices $W^{Q},\,W^{K},\,W^{V}\in\mathbb{R}^{d\times d}$, and $W_{O}\in\mathbb{R}^{d\times d}$. The columns of each projection matrix and rows of the outputs matrix can be split into $H$ equal parts, corresponding to the number of attention heads $W_{Q}^{\ell,j},\,W_{K}^{\ell,j},\,W_{V}^{\ell,j}\in\mathbb{R}^{d\times\tfrac{d}{H}}$, and $W_{O}^{\ell,j}\in\mathbb{R}^{\tfrac{d}{H}\times d},\,j=[1,H]$. Let $X^{\ell-1}\in\mathbb{R}^{N\times d}$ be the matrix of token representations at layer $\ell-1$. This allows the MHSA output as a sum of matrices, each induced by single attention head:
$a_{i}^{\ell}=\sum_{j=1}^{H}A^{\ell,j}\bigl{(}X^{\ell-1}\,W_{V}^{\ell,j}\bigr{)}\,W_{O}^{\ell,j}.$
After $L$ layers, the final hidden state $\mathrm{x}_{T}^{L}$ is mapped to vocabulary logits via an unembedding matrix $W_{U}\in\mathbb{R}^{|V|\times d}$, so that
$y=W_{U}\,\mathrm{x}_{T}^{L}\in\mathbb{R}^{|V|}$
and $\mathrm{softmax}(y)$ defines the next-token distribution. Thus we can view a transformer as a computational graph $G\colon\mathcal{X}\to\mathcal{Y}$, where token embeddings, MHSA sublayers, MLP sublayers, and the unembedding all communicate via the residual stream. We refer to Vaswani et al. (2017); Geva et al. (2023) for further details.
3.2 Knockout
If we think of a model as a computational graph $G$, then certain behaviors of the model can be attributed to specific subgraphs $G_{sub}\subseteq G$ Rai et al. (2024). Identifying such subgraphs provides a valuable mechanism for interpretability, enabling the isolation of individual components and their contributions to the model’s behavior. A knockout removes specific components to quantify their contribution to particular behaviors Nanda et al. (2023); Wang et al. (2022); Geva et al. (2023). Every activation in $G_{\text{sub}}$ can be replaced by either zero or reference mean:
$\mathrm{x}_{i}^{\ell}\;\longmapsto\;\tilde{\mathrm{x}}_{i}^{\ell}=\begin{cases}0\quad\text{or}\quad\mu_{i}^{\ell}&\text{if }(\ell,i)\in G_{\text{sub}},\\ \mathrm{x}_{i}^{\ell}&\text{otherwise}.\end{cases},$
then $G$ with those components “knocked out” is
$G_{\text{knock}}=G\left(\dots,\tilde{\mathrm{x}}_{i}^{\ell},\dots\right).$
By replacing the activations of these components, we eliminate the contributions of $G_{sub}$ while keeping other computations in $G$ fixed.
3.3 Activation patching
Activation patching is a technique to identify which activations in a model contribute to a particular output Wang et al. (2022); Meng et al. (2022); Hanna et al. (2023); Conmy et al. (2023); Stolfo et al. (2023). Instead of knocking out a component, we replace it with specific representations. This involves running the model on input A, then selectively replacing (patching) certain activations with those obtained from a paired input B:
$G_{\text{patch}}(A\hookleftarrow B)=G\bigl{(}\mathrm{x}^{1}_{i}(A),\dots,\mathrm{x}^{\ell}_{i}(B),\mathrm{x}^{\ell+1}_{i}(A),\dots\bigr{)}.$
By comparing the original output for input A with the altered output after patching, we can precisely measure how much the introduced activations from input B shift the model’s output.
3.4 Metric for measuring interpretation
To measure whether the model assumes the figurative or literal interpretation when given each prompt $s\in\{s_{a},s_{f},s_{l}\}$, we define two scores based on cumulative probabilities:
$F(s)\;=\;\sum_{c_{f}\in C_{f}}p(c_{f}\mid s);\;L(s)\;=\;\sum_{c_{l}\in C_{l}}p(c_{l}\mid s),$
where $C_{f}$ and $C_{l}$ are the sets of tokens corresponding to figurative and literal interpretation, respectively. After applying an intervention (knockout or activation patching), we quantify its causal effect:
$\Delta I(s)\;=\;I_{\mathrm{interv}}(s)\;-\;I_{\mathrm{origin}}(s),\,I\in\{F,L\},$
where $\Delta I<0$ means the intervention disrupts the corresponding interpretation, showing the targeted component is essential for that meaning. In contrast, $\Delta I=0$ indicates no effect and $\Delta I>0$ implies that the intervened component is not required (or may even been inhibited) that particular reading.
4 Localizing idiom meaning retrieval
4.1 Probing sublayers through knockout
To locate the mechanisms for retrieving figurative meanings from the model’s parameters, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$. By knocking out the activations across each layer, we assess how the probabilities of predicting specific candidate tokens ($\Delta F(s_{a}),\,\Delta L(s_{a})$) change. To narrow down the points to which we can ascribe a functional role for a specific computation component, we limit the intervention to the MLP and MHSA in each block.
(a) MLP
(b) MHSA
(c) MLP
(d) MHSA
Figure 3: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for contexts $s\in\{s_{a},s_{f},s_{l}\}$. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant difference between $\Delta F(s_{a})$ and the others (paired $t$-test, $p<0.05$). The difference at * marked layer is larger than the average difference across all layers.
Results Figure 3(a) shows by how much the model’s interpretation of the idiom changes when an MLP sublayer is knocked out. We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2): from $-0.57$ to $-0.25$. At the same time, we see an increase in the probability of the literal completions: $\Delta L(s_{a})$ rises above zero, from $0.37$ to $0.18$. And these shifts are significantly different from 0-2 layers ($p<0.05$). This indicates that early layer MLPs are diagnostic for reading out the figurative meaning while potentially suppressing the literal one. For later layers ($\geq$ 4), MLP sublayer knockout does not affect the model’s interpretation, which converges to zero.
Figure 3(b) shows the effect of attention knockout, which is similar to what we observed with MLP sublayer knockout. Specifically, knocking out MHSA leads to a drop in $\Delta F(s_{a})$ and increase in $\Delta L(s_{a})$ in the early layers (0-3). The most significant difference between $\Delta F(s_{a})$ and $\Delta L(s_{a})$ is in layer 1, where $\Delta F(s_{a})$ dropped by $-0.36$ and $\Delta L(s_{a})$ increased by $0.43$ ($p<0.05$). After layer 4, knocking out MHSA has little to no effect on interpretation, as the values converge toward zero.
Together, the two interventions track the 1 idiom’s figurative meaning retrieval: Early MHSA layers (0-3) gather and bind token-wise interactions, while early MLP layers (0-2) transform those bound representations into figurative meaning, while suppressing the literal meaning. This early layer idiom meaning retrieval aligns with the findings of Haviv et al. (2022), who argue that idiomatic information is accessed in the initial layers of LLMs during inference. After layer 4, $\Delta F(s_{a})$ and $\Delta L(s_{a})$ stabilize to zero, indicating that later sublayers are not adding new information and that the representations of the idiom span tokens are no longer used by any subsequent tokens.
4.2 Disentangling idiom specific processing from general semantic processing
To validate whether early layers in MHSA and MLP components are specifically essential for idiom processing, or generally important for semantic interpretation of a sentence, we conduct additional knockout experiments on literal and figurative paraphrases ($s_{l},\,s_{f}$). If general compositional semantic processing does not need to rely on the idiom retrieval components, we would expect that knocking out them when processing non-ambiguous sentences (i.e., paraphrases) would not disrupt the inference.
ResultsFigure 3(c) shows that the drop of knocking out early MLP sublayers is significantly smaller in layers 1 and 2 for the two paraphrases $\mathrm{min}$($\Delta L(s_{l})$, $\Delta F(s_{f})$)$=-0.11$ than for the idiomatic expression, $\mathrm{min}$($\Delta F(s_{a})$)$=-0.36$ ($p<0.05$). While, at layer 0, the knockout effect is large for all expressions, as in this case, the MLP block is processing the semantics of the sentence.
Figure 3(d) shows a similar effect for the early attention layers (0-2) where literal paraphrases are not strongly affected by knockout; $\mathrm{min}$($\Delta L(s_{l})$, $\Delta F(s_{f})$)$=-0.08$, $\mathrm{min}$($\Delta F(s_{a})$)$=-0.36$ ($p<0.05$). Taken together, these results reveal that early MLP and MHSA sublayers are essential for the model’s retrieval of an idiom’s figurative meaning, and that this process is notably different for idiomatic phrases from general semantic processing. Knocking out the early layers preserves the general semantic comprehension, but disrupts its ability to interpret idiom expressions. This dissociation implies that the model activates a specialized component that is necessary for idiom understanding, yet remains inactive during non-ambiguous literal semantic processing.
4.3 Identifying attention heads specific to retrieving figurative meanings
Next, we investigated which attention heads are specialized for retrieving an idiom’s figurative meaning, while suppressing its literal compositional interpretation. We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$ to identify the attention heads most impactful for interpretation.
(a) $\Delta F(s_{a})$
(b) $\Delta L(s_{a})$
Figure 4: Heatmaps of the (a) $\Delta F(s_{a})$ (b) $\Delta L(s_{a})$ when ablating individual attention heads at the idiom span. Idiomatic heads: Heads those are crucial for retrieving figurative meaning of idiom; $-\Delta F(s_{a})$ and $+\Delta L(s_{a})$.
If a head contributes to retrieving figurative meanings from idioms, we expect that knocking it out would cause the model to lose confidence in the figurative meaning of $s_{a}$, while boosting its literal interpretation. To operationalize this, we define idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$. Figure 4 shows that there is indeed a subset of attention heads that simultaneously exhibits large $-\Delta F(s_{a})$ (Figure 4(a)) and $+\Delta L(s_{a})$ (Figure 4(b)). To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples to identify, estimating 95% confidence intervals for each head’s contribution to the interpretation. We then rank all heads by, on the one hand, $-\Delta F(s_{a})$, and, on the other hand, $+\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders, obtaining our list of 20 idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ (selected attention heads listed in Appendix Table 1). Further analyses show that these heads process literal contexts differently compared both to random heads, and to heads that instead focus on compositional semantic processing $\mathcal{H}_{\mathrm{sem}}$ (Appendix B).
4.4 Causal role of MLP sublayers and idiomatic heads in figurative reading
To validate that the early MLP sublayers and the selected idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ are critical for the figurative reading, we patch their activations on $s_{a}$ into the same components when processing $s_{l}$. As a control, we patch randomly chosen MLP sublayers and heads. We quantify the effect on interpretation by $\Delta F(s_{l}\hookleftarrow s_{a})$ and $\Delta L(s_{l}\hookleftarrow s_{a})$.
ResultsPatching these components boosts the figurative interpretation, as quantified by $\Delta F(s_{l}\hookleftarrow s_{a})$ ($M=0.51,\,SD=0.59$) while suppressing the literal interpretation, $\Delta L(s_{l}\hookleftarrow s_{a})$ ($M=-0.92,\,SD=0.59$). In contrast, patching random components barely affects the figurative interpretation ($\Delta F(s_{l}\hookleftarrow s_{a})$: $M=0.07,\,SD=0.75$) even while suppressing the literal interpretation ($\Delta L(s_{l}\hookleftarrow s_{a})$: $M=-1.83,\,SD=1.39$). Effects on the experiment and the control are significantly different ($p<0.05$). These contrasting effects suggest that literal interpretation depends on scattered non-specialized heads, so perturbing random heads disrupts $\Delta L(s_{l})$. Conversely, idiomatic heads are specialized to enhancing the figurative meaning while suppressing literal interpretations, so that replacing their activations injects a figurative interpretation and amplifies $\Delta F(s_{l})$, while dropping $\Delta L(s_{l})$.
5 Tracing the interpretation flow
5.1 Locating interpretation representation in token position
After the retrieval process, the model encodes both literal and figurative meanings in its residual stream. We next investigate how these two representations are weighted or integrated in the final prediction. To locate the representation of the meanings, we calculate the mutual nearest-neighbor kernel alignment Huh et al. (2024); Cho et al. (2024) between sentence embeddings of meaning parts from $s_{f}$ and $s_{l}$ (i.e., exclude the ‘X (would)’ and ‘because X was so/a/the’) encoded by BGE M3 Chen et al. (2024) and hidden states from various token positions of idiom sentence $s_{a}$. This metric quantifies the similarity between representations, with higher values indicating greater semantic similarity.
(a) Last idiom token
(b) Subsequent token
(c) After subsequent token
(d) Last token of sentence
Figure 5: Kernel alignment between hidden states ($\mathrm{x}$) extracted from four different token positions of $s_{a}$ and semantic embeddings of paraphrases ($s_{f},\,s_{l}$).
ResultsWe plot the kernel alignment at four token positions in Figure 5. We observe that at the subsequent token (Fig. 5(b)), the kernel alignment between the idiom’s hidden states and its figurative meaning surpasses that of the literal meaning after layer 4. This suggests that by this point, the model has begun to favor the figurative interpretation ( 2 selective interpretation step) after retrieving figurative meaning.
[…]
Figure 6: Conceptual description of the activation patching experiments for tracing information flow (L = literal interpretation; F = figurative interpretation).
[…]
Figure 7: Layer-wise interpretation shift after patching in activations from $s_{f},\,s_{l},\,s_{a}^{*}$. The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ is larger than the average difference across all layers; these are statistically significant differences according to a paired $t$-test, $p<0.05$.
5.2 Analyzing layer-wise competing interpretation flow through activation patching
To dissociate the computational pathways of processing the figurative and literal interpretations, we employ activation patching at the subsequent token (i.e., ‘because’) across layers, the point where the figurative interpretation is most strongly encoded. We generate three conditions by pairing a source – a figurative context, a literal context, or a different idiom context ($s_{f},\,s_{l},\,s_{a}^{*}$), with a target sentence ($s_{a}$). Then, we patch the layer-wise residual activations at the ‘because’ position from the source condition into the target, see Figure 6 for illustration of expected effect. For each layer $\ell$, we measure the patching causal effect by $\Delta F$ and $\Delta L$. Figure 7 summarizes the shift of interpretation after patching. We additionally patch idiom representations into paraphrased sentences (Appendix 9), which support pathways of interpretations that we find.
ResultsPatching in figurative activations $s_{f}$ into $s_{a}$ removes the intermediate path literal flow but leaves the direct route intact (Figure 6(a)). Figure 7(a) shows that this slightly increases the $\Delta F(s_{a}\hookleftarrow s_{f})$ by up to $0.01$ for patching in mid-layers 4-10, but doesn’t affect the $\Delta L(s_{a}\hookleftarrow s_{f})$, which remains zero. This indicates that for the mid-layers at the ‘because’ token position, the model’s residual stream is the predominant channel for the figurative interpretation ( 3-1 figurative path). That is, we argue that there’s a different path for the literal interpretation ( 3-2 compositional semantic direct path) to convey its information to the final prediction, directly bypassing the intermediate path. This claim is supported by an attention-edge knockout experiment, where we prevent the last token from attending to the idiom tokens, and observe a drop in $\Delta L(s_{a})$ at layer 11 and 13 while $\Delta F(s_{a})$ remains near zero (see Appendix D).
We next patch the intermediate path by activations from the literal paraphrase $s_{l}$. This removes the intermediate path figurative flow (Figure 6(b)). Figure 7(b) shows a substantial increase in the $\Delta L(s_{a}\hookleftarrow s_{l})$, by $0.22$ in mid-layers (5-10). At the same time, $\Delta F(s_{a}\hookleftarrow s_{l})$ drops sharply at intermediate layers, which indicates that the figurative interpretation is severely disrupted.
By replacing idiom A’s intermediate representation with that of idiom B, we remove any information specific to idiom A that would have been transmitted through the intermediate pathway. Figure 7(c) shows that there is a drop of $\Delta F(s_{a}\hookleftarrow s_{a}^{*})$ by $-0.09$ in layer 7. On the other hand, the probability of the literal interpretation $\Delta L(s_{a}\hookleftarrow s_{a}^{*})$ does not change significantly. This implies that the literal path is intact, and there’s a bit of noise by adding the unrelated activations which cause a slight drop of $\Delta L(s_{a}\hookleftarrow s_{a}^{*})$ compared to $\Delta L(s_{a}\hookleftarrow s_{l})$. The magnitude of the drop in $\Delta F(s_{a}\hookleftarrow s_{a}^{*})$ closely matches the decrease after patching in the literal paraphrase $\Delta F(s_{a}\hookleftarrow s_{l})$, demonstrating that the figurative interpretation is predominantly routed through the intermediate pathway. In Appendix D, we further demonstrate that ablating the attention edge from the idiom span to its subsequent token causes $\Delta L(s_{a})$ to increase while $\Delta F(s_{a})$ decreases. This provides additional evidence that the figurative meaning is predominantly conveyed through this intermediate pathway. Meanwhile, when we patch activations from idiom B into the intermediate pathway of idiom A, we expect the intermediate representations to be overwritten by those of idiom B (Figure 6(c)). This would lead to increased prediction probabilities for both the figurative and literal interpretations associated with idiom B. As shown in Figure 7(d), both $\Delta F^{B}(s_{a}\hookleftarrow s_{a}^{*})$ and $\Delta L^{B}(s_{a}\hookleftarrow s_{a}^{*})$ increase, indicating that information supporting both interpretations is transmitted through the intermediate path. However, the magnitudes of these increases are not equal, suggesting that the pathway preferentially amplifies figurative interpretation over the other.
6 Conclusion
We present multiple steps of idiom processing in a transformer model (Figure 1), and identify distinct components and flows involved in the interpretation of figurative meaning. By isolating idiomatic expressions from the context, we gain insight into which components of the model are most sensitive to figurative interpretation compared to literal one. However, contextual information plays a crucial role in interpreting idioms, as it helps determine whether a given expression should be understood in its literal or figurative sense Holsinger (2013). Future research would explore how the model’s processing changes when disambiguating context is provided. By inserting disambiguating cues both before and after the idiom, we could examine the interactions between contextual clarification and idiom comprehension.
Moreover, an important next step is to extend our analysis to a larger number of idioms and develop a diagnostic framework that predicts, from the model’s internal activations alone, whether an expression’s figurative meaning is known. Concretely, one could observe the engagement of idiom-specialized attention heads and the activation of retrieval and binding pathways for every idiom in the set. If a given idiom fails to activate its dedicated heads, this would point to a breakdown in retrieval or binding. By looking at these activation patterns, we could quantify the model’s competence with respect to the idiom comprehension. Such an approach would not only deepen our understanding of how figurative meanings are encoded in transformer representations but also provide a useful guidance for pinpointing specific processing failures in idiom comprehension.
Limitations
While causal tracing methods have been widely used in recent work Geva et al. (2023); Dar et al. (2022); Meng et al. (2022); Heimersheim and Nanda ; Nanda et al. (2023), they only approximate the actual information stored in activations. Moreover, knocking out or replacing the activations can lead the model to out-of-distribution behaviors and cast doubt on the robustness of any interpretability claim. Our study is limited to LLama3.2-1B-base model, which is much smaller than the current state-of-the-art transformers, scaling these analyses to larger models remains an open challenge. Also, we simplify idiom processing by excluding context, even though context critically shapes idiomatic meaning. Altogether, these factors limit the generalizability of our results.
Ethics statement
One of the intended uses of the LLama-3.3-70B-instruct model is content generation, which aligns with our use of the model for generating the candidates for next token prediction task Grattafiori et al. (2024). The candidates generated by the language model may include biased or sensitive attributes (e.g., race, minority status), which reflects stereotypes that the language model already has (See the Appendix Table 3 for examples).
Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Starting Grant “Individualized Interaction in Discourse”, grant agreement No. 948878). Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102.
Appendix A Selected top-20 attention heads
We further extracted semantic heads $\mathcal{H}_{\mathrm{sem}}$, i.e., heads that show a positive $\Delta F(s_{a})$ and a negative $\Delta L(s_{a})$, which are processing ambiguous sentence in a literal way. Ablating them suppresses literal processing and restores figurative understanding. For comparison, we also randomly selected heads that are neither in $\mathcal{H}_{\mathrm{idiom}}$ nor $\mathcal{H}_{\mathrm{sem}}$.
Type top-$20$ attention heads (layer, head)
Idiomatic (0, 4), (0, 28), (1, 5), (2, 9), (0, 30), (0, 8), (0, 19), (2, 2), (1, 25), (1, 18), (4, 31), (1, 31), (0, 12), (0, 3), (2, 28), (2, 17), (2, 8), (2, 30), (1, 9), (4, 24)
Semantic (1, 3), (3, 28), (3, 30), (0, 10), (7, 16), (1, 20), (2, 4), (3, 12), (3, 14), (6, 27), (2, 19), (10, 11), (4, 26), (0, 18), (6, 31), (9, 22), (2, 5), (0, 9), (1, 19), (0, 11)
Random (9, 20), (10, 30), (2, 11), (2, 23), (1, 1), (9, 23), (13, 24), (8, 31), (3, 8), (13, 9), (10, 1), (5, 15), (7, 14), (7, 24), (4, 5), (14, 2), (9, 17), (4, 0), (14, 19), (6, 9)
Table 1: Top-$20$ attention head sets for each type.
Appendix B Activation divergence analysis between idiomatic and literal contexts
To examine how idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ process idioma differently compared to literal contexts, we focus on their activations over the idiom token span. Let =$\mathcal{I}=\{i_{1},\dots,i_{K}\}\subseteq\{1,\dots,T\}$ be the indices of the idiom span tokens. Note that the literal paraphrase $s_{l}$ preserves the same idiom token span, hence $|\mathcal{I}|$ is identical for both $s_{a}$ and $s_{l}$. For each head $h\in\mathcal{H}_{\mathrm{idiom}}$, we compute the averaged activation from idiom span value vectors:
$v_{a}\;=\;\frac{1}{\mathcal{I}}\sum_{i\in\mathcal{I}}\alpha^{(a)}_{h,i}\,V_{i},\;v_{l}\;=\;\frac{1}{\mathcal{I}}\sum_{i\in\mathcal{I}}\alpha^{(l)}_{h,i}\,V_{i},$
where $\alpha^{(o)}_{h,i}$ and $\alpha^{(l)}_{h,i}$ are the attention weights on token $i$ in $s_{a}$ and $s_{l}$, respectively, and $V_{i}$ is the corresponding value projection. These mean vectors capture the direction in the head’s value subspace over the idiom span under each context. We then measure their alignment via cosine similarity:
$\cos(v_{a},\,v_{l})=\frac{v_{a}\cdot v_{l}}{\|v_{a}\|\,\|v_{l}\|}.$
A low cosine similarity between $v_{a}$ and $v_{l}$ means that the head’s representational direction shifts substantially when processing the idiomatic sentence $s_{a}$ compared to its literal paraphrase $s_{l}$, revealing its specialization in processing ambiguous sentence and extracting its figurative meaning. In contrast, a high cosine similarity indicates that the head maintains a similar representation across both contexts, reflecting a more general semantic role. For comparison, we also compute cosine similarities for the semantic heads $\mathcal{H}_{\mathrm{sem}}$ and for a random set of heads $\mathcal{H}_{\mathrm{rand}}$.
For each sentence, we compute the cosine similarity $\cos(v_{a},v_{l})$ for every head in a given set of 20 attention heads, and then average these 20 values to obtain a single mean similarity (Figure 8). The cosine similarity across the dataset is shown in Figure 8. The $cos\_random$ ($M=0.87,\,SD=0.07$), $cos\_sem$ ($M=0.80,\,SD=0.09$) is significantly larger than $cos\_idiom$ ($M=0.76,\,SD=0.12$), ($p<0.05$).
Figure 8: Cosine similarities between the attention activations on idiomatic sentences ($s_{a}$) and their literal paraphrase ($s_{l}$) for the top-20 idiomatic heads, semantic heads, and random heads.
By confirming that the heads identified as specialized in retrieving the idioms’ figurative meaning indeed have divergent attention patterns when processing idioms compared to literal paraphrases, we strengthen the evidence that these heads are responsible for the unique processing required to interpret idiomatic expressions correctly.
Appendix C Layer-wise interpretation flow through activation patching
We additionally patch idiom’s because token activation into paraphrased sentences’ because token position. The conceptual details of the experiment is described as follow.
[…]
Figure 9: Description of the activation patching experiment for identifying interpretation flow. L represents literal interpretation; F represents figurative interpretation. (a) By replacing ‘because’ token $s_{f}$ activation with $s_{a}$, we expect literal interpretation to be added in the intermediate channel while keeping the direct route flow to be activated with naive figurative meaning. (b) By replacing ‘because’ token $s_{l}$ activation with $s_{a}$, we expect figurative interpretation to be added via intermediate path, while keeping both paths for literal interpretation.
[…]
Figure 10: Layer-wise interpretation shift $\Delta I(s_{\mathrm{x}}\hookleftarrow s_{a}),\,\mathrm{x}\in\{f,l\}$ after activation patching on the ‘because’ token across three patch directions. Y-axis: Mean values of $\Delta L(s_{\mathrm{x}}\hookleftarrow s_{a})$, $\Delta F(s_{\mathrm{x}}\hookleftarrow s_{a})$. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant shift between literal ($C_{l}$) and figurative ($C_{f}$) candidate probability (paired $t$-test, $p<0.05$).
As expected, by adding idiom’s interpretation via intermediate channel, since it’s adding literal interpretation, there’s a increase in $\Delta L(x_{f}\hookleftarrow s_{a})$ by $0.1$, whereas $\Delta F(x_{f}\hookleftarrow s_{a})$ drops by $-0.07$ across layers (Figure 10(a)). Whereas, when adding the idiom’s figurative interpretation into intermediate path, there’s a big increase in $\Delta F(s_{l}\hookleftarrow s_{a})$ by $0.26$ at layer 7 and decrease in $\Delta L(s_{l}\hookleftarrow s_{a})$ by $-0.17$. This implies that when there’s idiomatic representation passing the intermediate path, to make final prediction as close to figurative one, the model tries to amplify figurative meaning that is passing the intermediate path so that the total amount of activations for figurative interpretation to be same, which should be more than its literal one.
That is, injecting the idiom’s literal activation into the intermediate channel increases $\Delta L(x_{f}\hookleftarrow s_{a})$ by up to $0.1$ in the mid-layers (4-9). Conversely, injecting figurative activation raises $\Delta F(s_{l}\hookleftarrow s_{a})$ by $0.26$. This asymmetric increase indicates that when an idiomatic representation passes the intermediate path, the model actively amplifies figurative interpretation so it dominates the final prediction.
Appendix D Analyzing edge-level intermediate channel through attention knockout
To trace the flow of information, we perform fine-grained intervention on the MHSA sublayers, since they mediate communication between token positions, therefore, any critical information must pass through them Geva et al. (2023).
To locate the intermediate and direct path for each interpretation, we focus on the attention scores between idiom to subsequent token (i.e., because) for the intermediate path, and the idiom to last token of the sentence for the direct path findings. We investigate how this interpretation of flow is done internally by knockout the attention edges between target tokens. When critical attention edges for a certain interpretation of the idiom are blocked, it will result in severe degradation in probabilities of certain candidates. Therefore, we test whether critical information propagates between two hidden representations at a specific layer, by zeroing-out all the attention edges between them.
Formally, let $i,j\in\{1,\dots,T\}$ with $j\leq i$ be two positions in the sequence. To block token $i$ (query) from attending to token $j$ (key) at layer $\ell<L$, we overwrite the corresponding entries in the causal self-attention mask $M$ at layer $\ell+1$ across all heads $h\in\{1,\dots,H\}$:
$M_{i,\,j}^{\ell+1,\,h}\;=\;-\,\infty\quad\forall\,h\in\{1,\dots,H\}.$
(a) Idiom $\rightarrow$ Subsequent
(b) Idiom $\rightarrow$ Last
Figure 11: Layer-wise interpretation shift $\Delta I(s_{a})$ after knock out the attention edges between tokens. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant shift between literal ($C_{l}$) and figurative ($C_{f}$) candidate cumulative probabilities (paired $t$-test, $p<0.05$).
This restricts the query position from obtaining information from the key position, at that particular layer ($\ell+1$). For each layer, we block attention edges between (a) idiom span tokens $\rightarrow$ subsequent token (b) idiom span tokens $\rightarrow$ last token. If the specific path is important for a specific interpretation, then there should be significant drop of $I(s_{a})$ value. Figure 11 shows the result.
In Figure 11(a) at layer 5, it shows that there’s a drop in the $\Delta F(s_{a})$ by $-0.02$, while increase in $\Delta L(s_{a})$, by $0.02$. This shows that ablating attention from the idiom span to ‘because’ token disrupts the indirect path that would do disambiguation process where it suppresses the literal one. Without the idiom attending to the ‘because’ token, the model can no longer effectively disambiguate figurative meaning from literal meaning.
At layer 11 and 13, as in Figure 11(b), $\Delta L(s_{a})$ declines while $\Delta F(s_{a})$ rises. This divergence indicates that ablating the direct idiom to last token pathway prevents the model from passing the literal interpretation.
Ambiguous sentence Literal paraphrase Figurative paraphrase Literal candidates Figurative candidates
They will bend over backwards because they are so They will arch spine backwards because they are so They will make extra efforts because they are so flexible, used, strong, weak, relaxed, tight, uncomfortable, tall, stiff, short, scared, full, comfortable, tense, small, thin, angry, over, inf, surprised grateful, eager, proud, passionate, close, motivated, committed, keen, desperate, enthusiastic, invested, confident, glad, well, interested, sure, apprec, focused, attracted, dedicated
He bit off more than he can chew because he was so He took more food than he can swallow because he was so He took on more than he could handle because he was so hungry, greedy, happy, very, nervous, fam, attracted, r, thirsty, poor, star, pleased, starving, sad, delighted, hung, eng, drunk, gl, tempted eager, confident, anxious, desperate, optimistic, enthusiastic, passionate, sure, determined, ambitious, focused, driven, busy, good, full, young, keen, convinced, smart, strong
He would blow his own horn because he was a He would blow the musical instrument because he was a He would praise himself because he was a professional, musician, fan, p, wind, skilled, shepherd, trumpet, member, trump, pro, fl, virt, jazz, boy, bag, human, brass, flute, musical good, great, man, self, genius, proud, winner, god, true, hero, hard, legend, unique, better, brilliant, successful, clever, narciss, smart, perfection
It was out in left field because it was a It sat far in baseball’s area because it was a It was completely unrealistic because it was a baseball, league, minor, sport, strong, few, popular, difficult, download, home, football, smaller, tough, sports, right, significant, basketball, pitcher, deep, stadium fantasy, dream, very, completely, huge, one, product, total, movie, story, complete, perfect, romantic, fairy, totally, cartoon, fictional, massive, two, film
They were up the creek because they were a They were near the stream because they were a They were in trouble because they were a group, family, fishing, nom, part, hunting, water, tribe, pair, thirsty, party, people, river, fish, stream, traveling, farming, curious, type, peaceful new, long, threat, small, minority, bunch, few, mixed, very, day, single, large, bad, man, young, mess, tiny, mix, poor, relatively
Table 2: Examples of generated data.
Ambiguous sentence Literal paraphrase Figurative paraphrase Literal candidates Figurative candidates
They left him out in the cold because he was a They left him outside in the frost because he was a They abandoned him without support because he was a bad, stranger, little, bit, thief, trouble, poor, drunk, witch, dirty, dog, rebel, beg, danger, trait, s, nuisance, tiny, he, naughty child, foreign, threat, boy, political, baby, black, non, reminder, disabled, male, burden, cripp, product, minority, son, member, difficult, minor, different
They let him off the hook because he was a They removed him from fishing tackle because he was a They freed him from responsibility because he was a threat, bad, convicted, little, bit, danger, sexual, fish, ped, white, criminal, racist, bully, black, terrorist, thief, political, jerk, rap, big minor, child, foreign, slave, kid, good, mad, juvenile, victim, boy, stranger, young, youth, mere, student, fool, first, teenager, prisoner, friend
She was looking for a needle in a haystack because she was a She was searching for a neddle within a straw because she was a She was facing an impossible search because she was a needle, craft, seam, straw, tiny, hay, detective, master, haystack, witch, farmer, pro, camel, professional, busy, crazy, giant, chicken, cow, neat woman, victim, single, young, black, female, girl, slave, mother, stranger, new, first, non, private, white, novice, minority, prisoner, foreign, mom
Table 3: Biased data examples reflecting stereotypes embedded in the language model.

### v4

@@
-Tug-of-war between idiom’s figurative and literal meanings in LLMs
+Tug-of-war between idioms’ figurative and literal interpretations in LLMs
 Abstract
-Idioms present a unique challenge for language models due to their non-compositional figurative meanings, which often strongly diverge from the idiom’s literal interpretation.
-This duality requires a model to learn representing and deciding between the two meanings to interpret an idiom in a figurative sense, or literally.
-In this paper, we employ tools from mechanistic interpretability to trace how a large pretrained causal transformer (LLama3.2-1B-base) deals with this ambiguity.
-We localize three steps of idiom processing: First, the idiom’s figurative meaning is retrieved in early attention and MLP sublayers.
-We identify specific attention heads which boost the figurative meaning of the idiom while suppressing the idiom’s literal interpretation.
-The model subsequently represents the figurative representation through an intermediate path.
-Meanwhile, a parallel bypass route forwards literal interpretation, ensuring that a both reading remain available.
-Overall, our findings provide a mechanistic evidence for idiom comprehension in an autoregressive transformer.
-[^1] Code and dataset are available at https://github.com/sori424/idiom_processing under the MIT License.
-$\dagger$\$$\dagger$\$footnotetext: Co-corresponding authors.
+Idioms present a unique challenge for language models due to their non-compositional figurative interpretations, which often strongly diverge from the idiom’s literal interpretation.
+In this paper, we employ causal tracing to systematically analyze how pretrained causal transformers deal with this ambiguity.
+We localize three mechanisms: (i) Early sublayers and specific attention heads retrieve an idiom’s figurative interpretation, while suppressing its literal interpretation.
+(ii) When disambiguating context precedes the idiom, the model leverages it from the earliest layer and later layers refine the interpretation if the context conflicts with the retrieved interpretation.
+(iii) Then, selective, competing pathways carry both interpretations: an intermediate pathway prioritizes the figurative interpretation and a parallel direct route favors the literal interpretation, ensuring that both readings remain available.
+Our findings provide mechanistic evidence for idiom comprehension in autoregressive transformers [^1] Code and dataset are available at https://github.com/sori424/idiom_processing under the MIT License..
+[…]
 1 Introduction
-Figure 1: The figurative and literal interpretations are highlighted in the block and path.
-We find three main steps for idiom processing: 1 Idiom retrieval step: Early layers (i.e., 0-3) attention and MLP are actively retrieving the idiom’s figurative meaning while storing both figurative and literal interpretations in the residual stream.
+Figure 1: The figurative and literal interpretations are highlighted in the blocks and paths.
+We find three main steps for idiom processing: 1 Idiom retrieval step: Early layers (i.e., layers 0-3) attention and MLP are actively retrieving the idiom’s figurative interpretation while storing both figurative and literal interpretations in the residual stream.
 2 Selective interpretation step: At the token immediately following the idiom span, the model begins to encode a representation that favors the figurative interpretation over the literal one, starting from the middle layers.
-3 Interpretation routing: For final prediction, the model passes literal interpretation via both a direct compositional semantic path ( 3-2), as well as the intermediate pathway that prioritizes the figurative meaning ( 3-1 figurative path).
-Idioms challenge standard semantic composition because they are, by definition, multi-word sequences whose figurative meanings diverge from the composition of their literal parts Beck and Weber (2020).
-For instance, “kick the bucket” conveys the figurative meaning “to die”, yet its literal sense “physically kicking a container” remains semantically available.
-This ambiguity often leads large language models (LLMs) to misinterpret idioms, favoring their literal sense over the intended figurative one Kabra et al.
+3 Interpretation routing: For final prediction, the model passes literal interpretation via both a direct compositional semantic path ( 3-2), as well as the intermediate pathway that prioritizes the figurative interpretation ( 3-1 figurative path).
+Idioms pose a challenge to standard semantic composition because, as multi-word expressions, their figurative meanings don’t follow from the literal meanings Beck and Weber (2020).
+For instance, “kick the bucket” conveys the figurative interpretation “to die”, yet its literal sense “physically kicking a container” remains semantically available.
+Psycholinguistics has investigated how humans understand idioms, whether they access the figurative interpretation directly or first have to access the literal interpretation of the idiom and only in a later stage suppress that interpretation and access the figurative interpretation.
+These findings have led to various models describing human idiom processing Gibbs (1980); Cacciari and Glucksberg (1991); Bobrow and Bell (1973); Cacciari and Tabossi (1988).
+Yet for LLMs, it’s unclear how they process idioms, potentially limiting further improvements in their idiom comprehension Kabra et al.
 (2023); Liu et al.
@@
 (2024).
-Psycholinguistics has been investigated how humans understand idioms, whether they access the idiomatic meaning directly or first have to access the literal meaning of the idiom and only in a later stage suppress that meaning and access the figurative meaning.
-These findings have lead to various models describing human idiom processing Gibbs (1980); Cacciari and Glucksberg (1991); Bobrow and Bell (1973); Cacciari and Tabossi (1988).
-Yet for LLMs, it’s unclear how they process idioms, potentially limiting further improvements in their idiom comprehension.
 One explanation may lie in how the model transforms raw input token embeddings into richer representations Tian et al.
@@
 (2022).
-Within the autoregressive transformer, token representations are incrementally refined from the initial input embeddings to the final representation used for next token prediction Vig et al.
+Within the autoregressive transformer, token representations are incrementally refined through a series of residual layers, which consist of multi-head self-attention (MHSA) and MLP sublayers, each adding their outputs to update the representation Vig et al.
 (2020).
-This transformation is achieved through a series of residual layers, which consist of multi-head self-attention (MHSA) and MLP sublayers, each adding their outputs to update the representation.
-This multi-stage transformation suggests that idiom processing might follow a multi-step transformation process.
-Haviv et al.
-(2022) demonstrated that the early layers of transformer models store and retrieve idiomatic information, while the higher layers enhance model confidence.
-Meanwhile, Dankers et al.
-(2022) showed that during machine translation of idioms, the transformer model recognizes idiomatic expressions and processes them more as single lexical units.
-This involves grouping the idioms together using attention sublayers and reducing their interaction with surrounding context.
-However, both studies lack a detailed mapping of the specific model components and causal pathways underlying these idiom processing behaviors.
-In this paper, we first aim to identify how the idiom is interpreted by the transformer (i.e., LLama3.2-1B-base), where its figurative meaning is located, and how it is retrieved by the model.
-To this end, we pinpoint specific components of the model that have specialized in processing idioms by boosting their figurative meaning, while suppressing the compositional semantic literal interpretation.
-Having identified these, we proceed to investigating the information flow in the model, with the goal of identifying specific pathways through which the figurative meaning is passed forward to the final prediction, and whether this differs from the pathways of passing forward information about an idiom’s literal interpretation.
+This multi-stage transformation suggests that idiom processing proceeds via a sequential shift from compositional to non-compositional representations Haviv et al.
+(2023); Dankers et al.
+(2022).
+In this paper, we first aim to identify how the idiom is interpreted by a transformer Llama3.2-1B (we also run experiments on Llama3.1-8B, Qwen2.5-0.5B, Qwen2.5-7B), where its figurative interpretation is located, and how it is retrieved by the model compared to its literal counterpart.
+To this end, we pinpoint specific components of the model that have specialized in processing idioms by boosting their figurative interpretation, while suppressing the compositional semantic literal interpretation.
+Having identified these, we proceed to investigating the information flow in the model, with the goal of identifying specific pathways through which the figurative interpretation is passed forward to the final prediction, and whether this differs from the pathways of passing forward information about an idiom’s literal interpretation.
+Additionally, given that the context plays a key role in disambiguating the interpretation of an idiom Mi et al.
+(2025); Fazly et al.
+(2009); Dankers et al.
+(2022); Holsinger (2013), we investigate how context interacts with lexical information of idioms.
 We address these questions through knockout analyses Nanda et al.
@@
 (2022); Geva et al.
-(2023), where we separately knock out activations of each component (layer-wise MHSA and MLP, individual attention heads) in the model to observe their importance for retrieving idiom’s figurative meaning.
+(2023), where we separately ablate activations of each component (layer-wise MHSA and MLP, individual attention heads) in the model to observe their importance for retrieving idiom’s figurative interpretation, and contextual disambiguation.
 Furthermore, to trace the flow of these interpretations, we employ activation patching experiments Wang et al.
@@
 (2023), in which we ablate one or both interpretations from specific pathways within the model.
-Our main findings are summarized in Figure 1: we identify three main steps of idiom processing (1) retrieval of the figurative meaning of the idiom in the early layers, (2) selective encoding of the two interpretations at intermediate layers after the idiom span token position, and (3) two alternative paths for routing the idiom’s meaning to the final prediction, via a direct path for the literal meaning and an intermediate pathway that preferentially amplifies figurative interpretation over the other.
-In summary, our main contributions are:
-  • We propose a next token prediction task that distinguishes figurative vs. literal interpretations using interpretation-specific tokens (Section 2).
-  • We show that early layers contribute to retrieve the idiom’s figurative meaning, and that these are significantly different from general semantic processing (Section 4).
-  • We identify a selection mechanism at the token immediately following the idiom, where its residual activation shifts toward the figurative interpretation and then propagates through the intermediate layers (Section 5).
-  • We also uncover a bypass pathway that channels the literal interpretation directly into the final prediction, circumventing the intermediate route (Section 5.2).
+Figure 1 summarizes our main findings.
+Our contributions are:
+  • We find that early layers and idiom specific heads causally retrieve figurative interpretations, while suppressing the corresponding literal counterparts (Section 4).
+  • We identify a selection mechanism at the token immediately following the idiom, where an intermediate path carries a strong flow toward the figurative interpretation.
+In parallel, a bypass pathway preferentially carries literal information directly to the final prediction, circumventing the intermediate pathway (Section 5).
+  • We localize where context resolves idiomatic ambiguity, showing that early MLP/MHSA layers suppress figurative retrieval under literal context, while later MHSA integrates surrounding tokens to refine the context consistent interpretation (Section 6).
 2 Idiom interpretation task
-We are interested in teasing apart the language model’s processing of the literal vs.
-figurative interpretation of ambiguous idiom expressions.
-We therefore formalize a task and construct a dataset to investigate the interpretation shift in the model.
-In order to allow for effective probing, we embed the idiom into a sentence template, constructed such that the interpretation is revealed via next token prediction.
-The template [^2] X is instantiated with various pronouns (he/she/it/they); the word “would” is inserted only in some of the idioms to make the sentence sound more fluent; idioms are morphologically fit into the sentence; depending on sg/pl form of X, the second part of the template contained was or were.
-‘X (would) IDIOM because X was so/too/a/the’ can be applied to all of the selected idioms while preserving the idiom’s ambiguity.
-It elicits a causal completion, which we expect to differ between the literal and the figurative meaning of the idiom: for the figurative meaning of the idiom “kick the bucket”, we would expect completions that provide causes for the figurative meaning “to die”, e.g., old/sick, while for the literal meaning, we would expect completions that provide reasons for “kicking an object”, e.g.
-angry/mad.
-Figure 2 illustrates this process and selected examples can be found in Appendix Table 2.
-Figure 2: Pipeline for extracting figurative ($C_{f}$) and literal ($C_{l}$) candidate tokens for an idiomatic sentence.
-Since there is no single ground truth continuation that reveals either the figurative or the literal reading, we automatically construct sets of multiple tokens that are consistent with either the literal or the figurative meaning.
-The tokens in each set are mutually exclusive and selected based on their high likelihood of being associated with either the figurative or literal meaning of the idiom.
-In the following, we describe each step of the process of generating the dataset.
-1. Idiom Extraction: We select 245 idioms from a psycholinguistic paper Cronk et al. (1993), each of which received high literality ratings in a human rating study.
+We conduct our experiments in two setups.
+First, we investigate how language models process idiomatic expressions, which permit both figurative and literal interpretations.
+In the second setup, we introduce disambiguating context by adding preceding sentences.
+We formalize a next-token prediction task and construct a dataset.
+Idioms are embedded into a sentence template designed to reveal their interpretation through continuation ‘X (would) IDIOM because X was so/too/a/the’ [^2] Here, X is instantiated with pronouns and morphologically adapted; X is instantiated with various pronouns (he/she/it/they); the word “would” is inserted only in some of the idioms to make the sentence sound more fluent; idioms are morphologically fit into the sentence; depending on sg/pl form of X, the second part of the template contained was or were..
+This template preserves idiom ambiguity while eliciting causal completions (e.g., for ‘kick the bucket’, figurative completions include ‘old/sick’, while literal completions include ‘angry/mad’.) Since there is no single ground truth continuation, we automatically construct token sets consistent with either the figurative or literal interpretation.
+The tokens in each set are mutually exclusive and chosen for their high likelihood of association with the respective interpretation.
+Below, we outline the dataset construction process, and see Figure 9:
+1.
+Idiom Extraction: We select 245 idioms from a psycholinguistic paper Cronk et al.
+(1993), each of which received high literality (humans rated the literal meaning as plausible).
+This ensures ambiguity, which is necessary to evaluate the model’s change between literal and figurative interpretations.
 2.
@@
 […]
-.[$s_{a}$] He would kick the bucket because he was so
+$s_{a}$] He would kick the bucket because he was so
 […]
-.[$s_{f}$] He would die because he was so
+$s_{f}$] He would die because he was so
 […]
-.[$s_{l}$] He would kick the container because he was so
+$s_{l}$] He would kick the container because he was so
 3.
 Token Set Generation: To generate sets of next-word completions that are indicative of the figurative vs.
-the literal interpretation, we prompt the LLaMa3.3-70B model with disambiguated paraphrases ($s_{f}$, $s_{l}$) and obtain logits from the model with greedy decoding.
-Let $z^{(f)}$ and $z^{(l)}$ denote the logits corresponding to $s_{f}$ and $s_{l}$ as input, respectively.
-We then compute the element-wise difference between the two sets of logits as follows:
-$\Delta z_{v}=z_{v}^{(f)}-z_{v}^{(l)},$
-where $v\in V$ is the vocabulary size of the model (i.e., $V=$128,256).
-Since tokens that score high in one context tend to score low in the other, $\Delta z_{v}$ captures each token’s context-specific relevance.
-The 20 tokens that have highest $\Delta z_{k}$ are chosen as the set $C_{f}$ of tokens that indicate a figurative interpretation and the 20 tokens that have lowest $\Delta z_{k}$ values are chosen as the set $C_{l}$ that indicates a literal interpretation.
+the literal interpretation, we prompt the Llama3.3-70B-instruct model with paraphrases ($s_{f}$, $s_{l}$) and obtain logit distributions.
+Let $z^{(f)}$ and $z^{(l)}$ denote the logits corresponding to $s_{f}$ and $s_{l}$ as input, then we compute the element-wise difference between the two sets $\Delta z_{v}=z_{v}^{(f)}-z_{v}^{(l)}$, where $v\in V$ is the vocabulary size of the model.
+Since tokens that score high in one paraphrase tend to score low in the other, $\Delta z_{v}$ captures each token’s interpretation-specific relevance.
+We take the 20 tokens with the largest $\Delta z_{k}$ as the figurative candidate set $C_{f}$, and the 20 with the smallest $\Delta z_{k}$ as the literal candidate set $C_{l}$.
 4.
-Validation: To make sure that the Llama3.2-1B-base model has prior knowledge of the selected idioms, and the capability of discerning the context of the given sentence, we filter out the generated instances that satisfy the following two conditions:
-(1) Given $s_{a}$ as a prompt, the cumulative probabilities of predicting a figurative candidate as a next token must exceed that of the literal candidate.
-This metric quantifies that the model has the knowledge of the idiom:
-$\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{a})>\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{a}).$
-(2) Given $s_{f}$ and $s_{l}$ as prompt, the cumulative probabilities of predicting next token candidate matching the context exceeds that of the opposite context:
-$\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{f})>\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{f}),$
-$\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{l})>\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{l}).$
-110 out of the original set of 245 idioms passed these criteria. All following analyses are based on this subset of 110 validated idioms.
+Context Generation: Using each paraphrased sentence, we generate a preceding context sentence that disambiguates the interpretation of the subsequent idiom.
+The examples below illustrate the generated figurative context ($FC$) and literal context ($LC$).
+[…]
+.[FCFC] His breath rattled after every step. He would die because he was so
+[…]
+.[LCLC] The vending machine ate his coins. He would kick the container because he was so
 3 Methodology
 3.1 Information flow in transformers
-Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{T}]$ over a vocabulary $V$, each token is embedded as $\mathrm{x}_{i}^{0}=\mathrm{RoPE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{RoPE}$ applies rotary positional encoding to inject position $i$ Su et al.
+Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{T}]$ over a vocabulary $V$, each token is embedded as $\mathbf{x}_{i}^{0}=\mathrm{PE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{PE}$ denotes a positional encoding function that injects position $i$ into the token embedding Vaswani et al.
 (2017), and $d$ is the model’s hidden dimension.
@@
 where $W_{I}^{\ell}\in\mathbb{R}^{d_{ff}\times d}$ and $W_{F}^{\ell}\in\mathbb{R}^{d\times d_{ff}}$ are parameter matrices with inner-dimension $d_{ff}$ and $\sigma$ is a nonlinear activation function.
-MHSA SublayersEach attention layer’s output $\mathrm{a}^{\ell}_{i}$ aggregates information of multiple parallel attention heads $h$.
-The parameter matrices of each MHSA sublayer comprise three projection matrices $W^{Q},\,W^{K},\,W^{V}\in\mathbb{R}^{d\times d}$, and $W_{O}\in\mathbb{R}^{d\times d}$.
+MHSA SublayersEach attention layer’s output $a_{i}^{\ell}$ aggregates information of multiple parallel attention heads $h$.
+The parameter matrices of each MHSA sublayer comprise three projection matrices $W_{Q}^{\ell},\,W_{K}^{\ell},\,W_{V}^{\ell}\in\mathbb{R}^{d\times d}$, and $W_{O}^{\ell}\in\mathbb{R}^{d\times d}$.
 The columns of each projection matrix and rows of the outputs matrix can be split into $H$ equal parts, corresponding to the number of attention heads $W_{Q}^{\ell,j},\,W_{K}^{\ell,j},\,W_{V}^{\ell,j}\in\mathbb{R}^{d\times\tfrac{d}{H}}$, and $W_{O}^{\ell,j}\in\mathbb{R}^{\tfrac{d}{H}\times d},\,j=[1,H]$.
-Let $X^{\ell-1}\in\mathbb{R}^{N\times d}$ be the matrix of token representations at layer $\ell-1$.
-This allows the MHSA output as a sum of matrices, each induced by single attention head:
-$a_{i}^{\ell}=\sum_{j=1}^{H}A^{\ell,j}\bigl{(}X^{\ell-1}\,W_{V}^{\ell,j}\bigr{)}\,W_{O}^{\ell,j}.$
-After $L$ layers, the final hidden state $\mathrm{x}_{T}^{L}$ is mapped to vocabulary logits via an unembedding matrix $W_{U}\in\mathbb{R}^{|V|\times d}$, so that
-$y=W_{U}\,\mathrm{x}_{T}^{L}\in\mathbb{R}^{|V|}$
-and $\mathrm{softmax}(y)$ defines the next-token distribution.
+Let $X^{\ell-1}\in\mathbb{R}^{N\times d}$ be the matrix of all token representations at layer $\ell-1$, $A^{\ell,j}\in\mathbb{R}^{N\times N}$ encodes the weights computed by the $j$-th attention head at layer $\ell$, where it’s masked to a lower triangular matrix.
+The MHSA output is the sum over individual attention heads matrices:
+$a_{i}^{\ell}=\sum_{j=1}^{H}\big(A_{i}^{\ell,j}X^{\ell-1}W_{V}^{\ell,j}\big)W_{O}^{\ell,j}.$
+After $L$ layers, the final hidden state $\mathrm{x}_{T}^{L}$ is mapped to vocabulary logits via an unembedding matrix $W_{U}\in\mathbb{R}^{|V|\times d}$, so that $y=W_{U}\,\mathrm{x}_{T}^{L}\in\mathbb{R}^{|V|}$, and $\mathrm{softmax}(y)$ defines the next‐token distribution.
 Thus we can view a transformer as a computational graph $G\colon\mathcal{X}\to\mathcal{Y}$, where token embeddings, MHSA sublayers, MLP sublayers, and the unembedding all communicate via the residual stream.
@@
 (2023).
-Instead of knocking out a component, we replace it with specific representations.
-This involves running the model on input A, then selectively replacing (patching) certain activations with those obtained from a paired input B:
+This involves running the model on input A, then selectively patching certain activations with those obtained from a paired input B:
 $G_{\text{patch}}(A\hookleftarrow B)=G\bigl(\mathrm{x}^{1}_{i}(A),\dots,\mathrm{x}^{\ell}_{i}(B),\mathrm{x}^{\ell+1}_{i}(A),\dots\bigr).$
 By comparing the original output for input A with the altered output after patching, we can precisely measure how much the introduced activations from input B shift the model’s output.
-3.4 Metric for measuring interpretation
+3.4 Metric for measuring interpretation shift
 To measure whether the model assumes the figurative or literal interpretation when given each prompt $s\in\{s_{a},s_{f},s_{l}\}$, we define two scores based on cumulative probabilities:
-$F(s)\;=\;\sum_{c_{f}\in C_{f}}p(c_{f}\mid s);\;L(s)\;=\;\sum_{c_{l}\in C_{l}}p(c_{l}\mid s),$
-where $C_{f}$ and $C_{l}$ are the sets of tokens corresponding to figurative and literal interpretation, respectively.
-After applying an intervention (knockout or activation patching), we quantify its causal effect:
+$F(s)\;=\;\sum_{c_{f}\in C_{f}}p(c_{f}\mid s);\;L(s)\;=\;\sum_{c_{l}\in C_{l}}p(c_{l}\mid s).$
+After applying an intervention, we quantify its causal effect:
 $\Delta I(s)\;=\;I_{\mathrm{interv}}(s)\;-\;I_{\mathrm{origin}}(s),\,I\in\{F,L\},$
-where $\Delta I<0$ means the intervention disrupts the corresponding interpretation, showing the targeted component is essential for that meaning.
-In contrast, $\Delta I=0$ indicates no effect and $\Delta I>0$ implies that the intervened component is not required (or may even been inhibited) that particular reading.
-4 Localizing idiom meaning retrieval
+where $\Delta I<0$ means the intervention disrupts the corresponding interpretation, implying the component is necessary; $\Delta I=0$ shows no effect; $\Delta I>0$ implies the component is not required (or may even inhibit) that particular reading.
+4 Localizing idiom’s figurative interpretation retrieval
+We deliberately preserve idiomatic ambiguity by analyzing idioms without the context, so we can localize where, and through which components, the model shifts between figurative and literal interpretations.
 4.1 Probing sublayers through knockout
-To locate the mechanisms for retrieving figurative meanings from the model’s parameters, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$.
+To locate the mechanisms for retrieving figurative interpretations from the model’s parameters, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$.
 By knocking out the activations across each layer, we assess how the probabilities of predicting specific candidate tokens ($\Delta F(s_{a}),\,\Delta L(s_{a})$) change.
@@
 (d) MHSA
-Figure 3: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for contexts $s\in\{s_{a},s_{f},s_{l}\}$.
+Figure 2: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for sentences $s\in\{s_{a},s_{f},s_{l}\}$.
 Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals.
@@
 Results Figure 2(a) shows by how much the model’s interpretation of the idiom changes when an MLP sublayer is knocked out.
-We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2): from $-0.57$ to $-0.25$.
-At the same time, we see an increase in the probability of the literal completions: $\Delta L(s_{a})$ rises above zero, from $0.37$ to $0.18$.
-And these shifts are significantly different from 0-2 layers ($p<0.05$).
-This indicates that early layer MLPs are diagnostic for reading out the figurative meaning while potentially suppressing the literal one.
-For later layers ($\geq$ 4), MLP sublayer knockout does not affect the model’s interpretation, which converges to zero.
-Figure 3(b) shows the effect of attention knockout, which is similar to what we observed with MLP sublayer knockout.
-Specifically, knocking out MHSA leads to a drop in $\Delta F(s_{a})$ and increase in $\Delta L(s_{a})$ in the early layers (0-3).
-The most significant difference between $\Delta F(s_{a})$ and $\Delta L(s_{a})$ is in layer 1, where $\Delta F(s_{a})$ dropped by $-0.36$ and $\Delta L(s_{a})$ increased by $0.43$ ($p<0.05$).
+We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2).
+At the same time, we see an increase in the probability of the literal completions: $\Delta L(s_{a})$ rises above zero.
+And these shifts are significantly different in 0-3 layers ($p<0.05$).
+This indicates that early layer MLPs are causally important for reading out the figurative interpretation while potentially suppressing the literal one.
+For later layers ($\geq$ 4), MLP sublayer knockout does not affect the model’s interpretation, where the corresponding patching effect converges to zero.
+Figure 2(b) shows the effect of MHSA knockout, which is similar to what we observed with MLP sublayer knockout.
+Specifically, knocking out early layers MHSA (0-3) leads to a drop in $\Delta F(s_{a})$ and increase in $\Delta L(s_{a})$.
+The most significant difference between $\Delta F(s_{a})$ and $\Delta L(s_{a})$ is in layer 1, where $\Delta F(s_{a})$ dropped by 0.30-0.30 and $\Delta L(s_{a})$ increased by $0.16$ ($p<0.05$).
 After layer 4, knocking out MHSA has little to no effect on interpretation, as the values converge toward zero.
-Together, the two interventions track the 1 idiom’s figurative meaning retrieval: Early MHSA layers (0-3) gather and bind token-wise interactions, while early MLP layers (0-2) transform those bound representations into figurative meaning, while suppressing the literal meaning.
-This early layer idiom meaning retrieval aligns with the findings of Haviv et al.
+Together, the two interventions track the 1 idiom’s figurative interpretation retrieval: Early MHSA layers (0-3) gather and bind token-wise interactions, while early MLP layers (0-3) transform those bound representations into a figurative interpretation, while suppressing the literal one.
+This early layer idiom interpretation retrieval aligns with the findings of Haviv et al.
 (2023), who argue that idiomatic information is accessed in the initial layers of LLMs during inference.
-After layer 4, $\Delta F(s_{a})$ and $\Delta L(s_{a})$ stabilize to zero, indicating that later sublayers are not adding new information and that the representations of the idiom span tokens are no longer used by any subsequent tokens.
-4.2 Disentangling idiom specific processing from general semantic processing
-To validate whether early layers in MHSA and MLP components are specifically essential for idiom processing, or generally important for semantic interpretation of a sentence, we conduct additional knockout experiments on literal and figurative paraphrases ($s_{l},\,s_{f}$).
-If general compositional semantic processing does not need to rely on the idiom retrieval components, we would expect that knocking out them when processing non-ambiguous sentences (i.e., paraphrases) would not disrupt the inference.
-ResultsFigure 3(c) shows that the drop of knocking out early MLP sublayers is significantly smaller in layers 1 and 2 for the two paraphrases $\mathrm{min}$($\Delta L(s_{l})$, $\Delta F(s_{f})$)$=-0.11$ than for the idiomatic expression, $\mathrm{min}$($\Delta F(s_{a})$)$=-0.36$ ($p<0.05$).
+After layer 4, $\Delta F(s_{a})$ and $\Delta L(s_{a})$ stabilize to zero, indicating that later sublayers are not adding new information and that the representations of the interpretations at the idiom span tokens are no longer used by any subsequent tokens.
+Moreover, these components are specifically essential for idiom processing, not generally important for semantic interpretation of a sentence.
+Figure 2(c) shows that the drop from knocking out early MLP sublayers is significantly smaller in layer 1 for the two paraphrases ($\Delta L(s_{l})$, $\Delta F(s_{f})$) than for the idiomatic expression, ($\Delta F(s_{a})$), ($p<0.05$).
 While, at layer 0, the knockout effect is large for all expressions, as in this case, the MLP block is processing the semantics of the sentence.
-Figure 3(d) shows a similar effect for the early attention layers (0-2) where literal paraphrases are not strongly affected by knockout; $\mathrm{min}$($\Delta L(s_{l})$, $\Delta F(s_{f})$)$=-0.08$, $\mathrm{min}$($\Delta F(s_{a})$)$=-0.36$ ($p<0.05$).
-Taken together, these results reveal that early MLP and MHSA sublayers are essential for the model’s retrieval of an idiom’s figurative meaning, and that this process is notably different for idiomatic phrases from general semantic processing.
-Knocking out the early layers preserves the general semantic comprehension, but disrupts its ability to interpret idiom expressions.
-This dissociation implies that the model activates a specialized component that is necessary for idiom understanding, yet remains inactive during non-ambiguous literal semantic processing.
-4.3 Identifying attention heads specific to retrieving figurative meanings
-Next, we investigated which attention heads are specialized for retrieving an idiom’s figurative meaning, while suppressing its literal compositional interpretation.
-We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$ to identify the attention heads most impactful for interpretation.
+Figure 2(d) shows a similar effect for the early MHSA layers (0-2) where literal paraphrases are not strongly affected by knockout ($\Delta L(s_{l})$, $\Delta F(s_{f})$) compared to idiomatic expression ($\Delta F(s_{a})$), ($p<0.05$).
+The crucial role of figurative interpretation retrieval in early layers is consistent across different models (Appendix D).
+4.2 Identifying attention heads specific to retrieving figurative interpretations
+Next, we investigated which attention heads are specialized for retrieving an idiom’s figurative interpretation.
+We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$.
 (a) $\Delta F(s_{a})$
@@
 Figure 3: Heatmaps of the (a) $\Delta F(s_{a})$ (b) $\Delta L(s_{a})$ when ablating individual attention heads at the idiom span.
-Idiomatic heads: Heads those are crucial for retrieving figurative meaning of idiom; $-\Delta F(s_{a})$ and $+\Delta L(s_{a})$.
-If a head contributes to retrieving figurative meanings from idioms, we expect that knocking it out would cause the model to lose confidence in the figurative meaning of $s_{a}$, while boosting its literal interpretation.
+Idiomatic heads: Heads those are crucial for retrieving the figurative interpretation of idiom; $\Delta F(s_{a})$ and $\Delta L(s_{a})$.
 To operationalize this, we define idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$.
-Figure 4 shows that there is indeed a subset of attention heads that simultaneously exhibits large $-\Delta F(s_{a})$ (Figure 4(a)) and $+\Delta L(s_{a})$ (Figure 4(b)).
-To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples to identify, estimating 95% confidence intervals for each head’s contribution to the interpretation.
-We then rank all heads by, on the one hand, $-\Delta F(s_{a})$, and, on the other hand, $+\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders, obtaining our list of 20 idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ (selected attention heads listed in Appendix Table 1).
-Further analyses show that these heads process literal contexts differently compared both to random heads, and to heads that instead focus on compositional semantic processing $\mathcal{H}_{\mathrm{sem}}$ (Appendix B).
+Figure 3 shows that there is a subset of attention heads that simultaneously exhibits large $\Delta F(s_{a})$ (Figure 3(a)) and $\Delta L(s_{a})$ (Figure 3(b)).
+To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples for each head’s contribution to the interpretation.
+We then rank all heads by, on the one hand, $\Delta F(s_{a})$, and, on the other hand, $\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders (selected attention heads listed in Appendix Table 1).
 4.3 Causal role of MLP sublayers and idiomatic heads in figurative reading
-To validate that the early MLP sublayers and the selected idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ are critical for the figurative reading, we patch their activations on $s_{a}$ into the same components when processing $s_{l}$.
+To validate that the early MLP sublayers and the selected idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ are critical for the figurative reading, we patch the activations on $s_{a}$ (at idiom span) into the corresponding components when processing $s_{l}$.
 As a control, we patch randomly chosen MLP sublayers and heads.
-We quantify the effect on interpretation by $\Delta F(s_{l}\hookleftarrow s_{a})$ and $\Delta L(s_{l}\hookleftarrow s_{a})$.
-ResultsPatching these components boosts the figurative interpretation, as quantified by $\Delta F(s_{l}\hookleftarrow s_{a})$ ($M=0.51,\,SD=0.59$) while suppressing the literal interpretation, $\Delta L(s_{l}\hookleftarrow s_{a})$ ($M=-0.92,\,SD=0.59$).
-In contrast, patching random components barely affects the figurative interpretation ($\Delta F(s_{l}\hookleftarrow s_{a})$: $M=0.07,\,SD=0.75$) even while suppressing the literal interpretation ($\Delta L(s_{l}\hookleftarrow s_{a})$: $M=-1.83,\,SD=1.39$).
+ResultsPatching these components boosts the figurative interpretation, as quantified by $\Delta F(s_{l}\hookleftarrow s_{a})$ (M=0.17,SD=0.49M=0.17,\,SD=0.49) while suppressing the literal interpretation, $\Delta L(s_{l}\hookleftarrow s_{a})$ (M=0.29,SD=0.58M=-0.29,\,SD=0.58).
+In contrast, patching random components barely affects the figurative interpretation ($\Delta F(s_{l}\hookleftarrow s_{a})$: M=0.008,SD=0.09M=-0.008,\,SD=0.09) as well as literal one ($\Delta L(s_{l}\hookleftarrow s_{a})$: M=0.005,SD=0.11M=-0.005,\,SD=0.11).
 Effects on the experiment and the control are significantly different ($p<0.05$).
-These contrasting effects suggest that literal interpretation depends on scattered non-specialized heads, so perturbing random heads disrupts $\Delta L(s_{l})$.
-Conversely, idiomatic heads are specialized to enhancing the figurative meaning while suppressing literal interpretations, so that replacing their activations injects a figurative interpretation and amplifies $\Delta F(s_{l})$, while dropping $\Delta L(s_{l})$.
+These contrasting effects suggest that idiomatic components are specialized to enhancing the figurative interpretation while suppressing literal interpretations, so that replacing their activations injects a figurative interpretation and amplifies $\Delta F(s_{l})$, while dropping $\Delta L(s_{l})$.
 5 Tracing the interpretation flow
-5.1 Locating interpretation representation in token position
-After the retrieval process, the model encodes both literal and figurative meanings in its residual stream.
+5.1 Locating idiom interpretation across token positions
+After the retrieval process, the model encodes both literal and figurative interpretations in its residual stream.
 We next investigate how these two representations are weighted or integrated in the final prediction.
-To locate the representation of the meanings, we calculate the mutual nearest-neighbor kernel alignment Huh et al.
+To locate the representation of the interpretations, we calculate the mutual nearest-neighbor kernel alignment Huh et al.
 (2024); Cho et al.
-(2024) between sentence embeddings of meaning parts from $s_{f}$ and $s_{l}$ (i.e., exclude the ‘X (would)’ and ‘because X was so/a/the’) encoded by BGE M3 Chen et al.
+(2025) between sentence embeddings of a corresponding idiom token span from $s_{f}$ and $s_{l}$ encoded by BGE M3 Chen et al.
 (2024) and hidden states from various token positions of idiom sentence $s_{a}$.
@@
 (a) Last idiom token
-(b) Subsequent token
-(c) After subsequent token
+(b) ‘because’ token
+(c) After ‘because’ token
 (d) Last token of sentence
 Figure 4: Kernel alignment between hidden states ($\mathrm{x}$) extracted from four different token positions of $s_{a}$ and semantic embeddings of paraphrases ($s_{f},\,s_{l}$).
-ResultsWe plot the kernel alignment at four token positions in Figure 5.
-We observe that at the subsequent token (Fig.
-5(b)), the kernel alignment between the idiom’s hidden states and its figurative meaning surpasses that of the literal meaning after layer 4.
-This suggests that by this point, the model has begun to favor the figurative interpretation ( 2 selective interpretation step) after retrieving figurative meaning.
+ResultsWe observe that at the ‘because’ token (Figure 4(b)), the kernel alignment between the idiom’s hidden states and its figurative interpretation surpasses that of the literal interpretation after layer 4.
+This suggests that by this point, the model has begun to favor the figurative interpretation ( 2 selective interpretation step).
+Based on the results, we hypothesize an intermediate pathway where token position ‘because’ plays a pivotal role in steering the idiom toward a figurative interpretation.
+In parallel, to keep the alternative interpretation available at the prediction (as in Figure 4(d)), we posit a direct route that favors the literal interpretation by more strongly preserving compositional semantic information.
+5.2 Analyzing layer-wise competing interpretation flow through activation patching
+To measure the information flow along the intermediate pathway (through the subsequent token position ‘because’), we patch the activation $s_{a}$ at the ‘because’ token with $s_{f}$ or $s_{l}$ (Figures 5(a), 5(b)).
+On the other hand, patching the activation $s_{a}$ at the idiom token span with an alternative activation ($s_{f}$ or $s_{l}$) and then re-patching attention at the ‘because’ token with $s_{a}$ allows us to isolate the information that flows via the direct route (Figures 5(c), 5(d)).
+[…]
 […]
@@
 […]
-Figure 7: Layer-wise interpretation shift after patching in activations from $s_{f},\,s_{l},\,s_{a}^{*}$.
-The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ is larger than the average difference across all layers; these are statistically significant differences according to a paired $t$-test, $p<0.05$.
-5.2 Analyzing layer-wise competing interpretation flow through activation patching
-To dissociate the computational pathways of processing the figurative and literal interpretations, we employ activation patching at the subsequent token (i.e., ‘because’) across layers, the point where the figurative interpretation is most strongly encoded.
-We generate three conditions by pairing a source – a figurative context, a literal context, or a different idiom context ($s_{f},\,s_{l},\,s_{a}^{*}$), with a target sentence ($s_{a}$).
-Then, we patch the layer-wise residual activations at the ‘because’ position from the source condition into the target, see Figure 6 for illustration of expected effect.
-For each layer $\ell$, we measure the patching causal effect by $\Delta F$ and $\Delta L$.
-Figure 7 summarizes the shift of interpretation after patching.
-We additionally patch idiom representations into paraphrased sentences (Appendix 9), which support pathways of interpretations that we find.
-ResultsPatching in figurative activations $s_{f}$ into $s_{a}$ removes the intermediate path literal flow but leaves the direct route intact (Figure 6(a)).
-Figure 7(a) shows that this slightly increases the $\Delta F(s_{a}\hookleftarrow s_{f})$ by up to $0.01$ for patching in mid-layers 4-10, but doesn’t affect the $\Delta L(s_{a}\hookleftarrow s_{f})$, which remains zero.
-This indicates that for the mid-layers at the ‘because’ token position, the model’s residual stream is the predominant channel for the figurative interpretation ( 3-1 figurative path).
-That is, we argue that there’s a different path for the literal interpretation ( 3-2 compositional semantic direct path) to convey its information to the final prediction, directly bypassing the intermediate path.
-This claim is supported by an attention-edge knockout experiment, where we prevent the last token from attending to the idiom tokens, and observe a drop in $\Delta L(s_{a})$ at layer 11 and 13 while $\Delta F(s_{a})$ remains near zero (see Appendix D).
-We next patch the intermediate path by activations from the literal paraphrase $s_{l}$.
-This removes the intermediate path figurative flow (Figure 6(b)).
-Figure 7(b) shows a substantial increase in the $\Delta L(s_{a}\hookleftarrow s_{l})$, by $0.22$ in mid-layers (5-10).
-At the same time, $\Delta F(s_{a}\hookleftarrow s_{l})$ drops sharply at intermediate layers, which indicates that the figurative interpretation is severely disrupted.
-By replacing idiom A’s intermediate representation with that of idiom B, we remove any information specific to idiom A that would have been transmitted through the intermediate pathway.
-Figure 7(c) shows that there is a drop of $\Delta F(s_{a}\hookleftarrow s_{a}^{*})$ by $-0.09$ in layer 7.
-On the other hand, the probability of the literal interpretation $\Delta L(s_{a}\hookleftarrow s_{a}^{*})$ does not change significantly.
-This implies that the literal path is intact, and there’s a bit of noise by adding the unrelated activations which cause a slight drop of $\Delta L(s_{a}\hookleftarrow s_{a}^{*})$ compared to $\Delta L(s_{a}\hookleftarrow s_{l})$.
-The magnitude of the drop in $\Delta F(s_{a}\hookleftarrow s_{a}^{*})$ closely matches the decrease after patching in the literal paraphrase $\Delta F(s_{a}\hookleftarrow s_{l})$, demonstrating that the figurative interpretation is predominantly routed through the intermediate pathway.
-In Appendix D, we further demonstrate that ablating the attention edge from the idiom span to its subsequent token causes $\Delta L(s_{a})$ to increase while $\Delta F(s_{a})$ decreases.
-This provides additional evidence that the figurative meaning is predominantly conveyed through this intermediate pathway.
-Meanwhile, when we patch activations from idiom B into the intermediate pathway of idiom A, we expect the intermediate representations to be overwritten by those of idiom B (Figure 6(c)).
-This would lead to increased prediction probabilities for both the figurative and literal interpretations associated with idiom B.
-As shown in Figure 7(d), both $\Delta F^{B}(s_{a}\hookleftarrow s_{a}^{*})$ and $\Delta L^{B}(s_{a}\hookleftarrow s_{a}^{*})$ increase, indicating that information supporting both interpretations is transmitted through the intermediate path.
-However, the magnitudes of these increases are not equal, suggesting that the pathway preferentially amplifies figurative interpretation over the other.
+Figure 6: Layer-wise interpretation shift after patching in activations.
+The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
+ResultsFigure 6(a) shows that the patching as illustrated in Figure 5(a), increases the $\Delta F(s_{a}\hookleftarrow s_{f})$ in mid-layers 4-10 (max: $0.17$), but has a relatively small effect on $\Delta L(s_{a}\hookleftarrow s_{f})$, which shows some decrease (min: 0.09-0.09).
+Patching in $s_{l}$ (as in Fig.
+5(b)) removes the intermediate path figurative flow, which leads to an increase in the $\Delta L(s_{a}\hookleftarrow s_{l})$ in mid-layers (6-10) as shown in Figure 6(b) (max: $0.13$).
+At the same time, $\Delta F(s_{a}\hookleftarrow s_{l})$ drops symmetrically in the same layers (min: 0.10-0.10).
+This indicates that the figurative interpretation is suppressed in favor of the literal interpretation.
+We interpret this as showing that a figurative injection doesn’t lead to suppression of the literal interpretation (an asymmetry between increases and decreases).
+By contrast, a literal injection must first suppress the figurative interpretation in that subspace to have an effect, producing a more symmetric push–pull tradeoff.
+This indicates that for the mid-layers at the ‘because’ token position, the model’s residual stream is the predominant route for the figurative interpretation ( 3-1 figurative path).
+Meanwhile, Figure 4(d) indicates that literal interpretation remains competitive with the figurative one.
+We assume that there’s a different path that favors the semantically literal interpretation ( 3-2 compositional semantic direct path) to convey its information to the final prediction, directly bypassing the intermediate path.
+To probe this, we replace the idiom sentence’s idiom token span activation with the activation from a paraphrased sentence, and re-patch the ‘because’ token with the idiom sentence’s original activation (as shown in Figures 5(c), 5(d)).
+This preserves the intermediate pathway and isolates the contribution of the direct route.
+Figure 6(c) shows that removing direct path’s literal interpretation produces a large drop in $\Delta L$ (min: 0.25-0.25) and rise in $\Delta F$ (max: $0.20$).
+Conversely, removing the direct path’s figurative interpretation yields only a modest drop in $\Delta F$ (min: 0.10-0.10) and a small increase in $\Delta L$ (max: $0.10$).
+These results indicate that the direct path predominantly conveys literal information.
+The pattern is consistent across different models (Appendix E).
+6 Localizing idiom disambiguation in context
+To assess the role of context in idiom disambiguation, we’re now looking at the setup with context.
+We retain an instance only when the model assigns higher cumulative probability to the context-consistent candidate than to the alternative (i.e., $F(s_{a}|FC)>F(s_{a}|LC)$, $L(s_{a}|LC)>L(s_{a}|FC)$).
+6.1 Probing sublayers through knockout with figurative vs. literal context
+We conduct a layer-wise knockout experiment to assess how MLP and MHSA components contribute to idiom processing when the same idiomatic expressions are presented in different contexts.
+(a) MLP FC
+(b) MLP LC
+(c) MHSA FC
+(d) MHSA LC
+Figure 7: Sublayer-wise interpretation shift $\Delta I$ after knockout activations at idiom span, for contexts FCFC: Figurative Context; LCLC: Literal Context.
+ResultsFigure 7(a) shows that for idioms in a figurative context, knocking out the MLP in the early layers (0–3) results in a significant drop in $\Delta F$ (min: 0.62-0.62) and increase in $\Delta L$ (max: $0.66$), consistent with the pattern observed for idioms without context (Figure 2(a)).
+This indicates that these layers are critical for retrieving the figurative interpretation independently of context.
+This result is consistent across different models (Appendix F).
+For idioms in a literal context (7(b)), ablating the MLP layer 1 component leads to a significant drop in $\Delta F$ (min: 0.33-0.33) and an increase in $\Delta L$ (max: $0.09$), indicating that the figurative interpretation is still being retrieved.
+The effect persists but is small compared with the figurative context condition, where early layers at the idiom token appear to encode context that suppresses figurative retrieval while biasing toward the literal interpretation.
+A similar pattern is observed in the MHSA layers.
+As shown in Figure 7(c), under figurative context, ablating early layers (0–3) produces a pronounced decrease in $\Delta F$ (min: 0.33-0.33) and an increase in $\Delta L$ (max: $0.48$).
+Under literal context, ablating the earliest layers (0–1) yields a drop in $\Delta F$ (min: 0.32-0.32), whereas only modest changes for $\Delta L$ (max: $0.16$) relative to the figurative context condition.
+That is, the literal context suppresses retrieval of the figurative interpretation from the very earliest layers.
+Moreover, ablating the layers (3–7) under literal context flips the pattern, where $\Delta F$ is increasing while $\Delta L$ is decreasing with a significant difference (Figure 7(d)).
+This indicates that, after the initial retrieval of figurative interpretation, the literal interpretation, which is disambiguated from context, is refined through mainly MHSA mediated integration with surrounding tokens.
+In sum, the results point to a two stage context disambiguation: (a) early layers retrieve figurative evidence while beginning to encode context, followed by (b) mid-layers, MHSA-driven contextual disambiguation that resolves conflicts in favor of the context-appropriate interpretation.
+6.2 Probing MHSA-mediated contextual disambiguation through KVQ patching
+To examine in detail how MHSA mediates contextual disambiguation in early-mid layers, we perform KVQ patching, in which MHSA’s key–value (KV) activations are replaced while selectively preserving idiom tokens’ query (Q) activations across layers, thereby steering the model’s attention toward alternative contexts.
+That is, we redirect idiom token span’s attention from a figurative-context to the literal context, and vice versa.
+If the model’s final prediction changes from figurative to literal (or vice versa) after patching, it demonstrates that the MHSA mechanism in those specific layers is responsible for using the context to steer the disambiguation process.
+[…]
+[…]
+[…]
+ResultsAs shown in Figure 8, patching in KVs with alternative context yields a shift in interpretation of idiom.
+Under FC, effects appears already at layer 0, indicating immediate context sensitivity, with additional peaks at layers 2-3 and 6.
+Under LC, the probability of the context matched interpretation drops across layers 2–6, while the alternative interpretation rises.
+Taken together, these results suggest that the idiom tokens exploit contextual evidence primarily in layers 2-6, after the figurative interpretation is first retrieved in layers 0-1, with layer 0 playing a crucial early role under FC.
+Later layers tend to consolidate rather than redirect the interpretation.
 7 Conclusion
-We present multiple steps of idiom processing in a transformer model (Figure 1), and identify distinct components and flows involved in the interpretation of figurative meaning.
-By isolating idiomatic expressions from the context, we gain insight into which components of the model are most sensitive to figurative interpretation compared to literal one.
-However, contextual information plays a crucial role in interpreting idioms, as it helps determine whether a given expression should be understood in its literal or figurative sense Holsinger (2013).
-Future research would explore how the model’s processing changes when disambiguating context is provided.
-By inserting disambiguating cues both before and after the idiom, we could examine the interactions between contextual clarification and idiom comprehension.
-Moreover, an important next step is to extend our analysis to a larger number of idioms and develop a diagnostic framework that predicts, from the model’s internal activations alone, whether an expression’s figurative meaning is known.
-Concretely, one could observe the engagement of idiom-specialized attention heads and the activation of retrieval and binding pathways for every idiom in the set.
-If a given idiom fails to activate its dedicated heads, this would point to a breakdown in retrieval or binding.
-By looking at these activation patterns, we could quantify the model’s competence with respect to the idiom comprehension.
-Such an approach would not only deepen our understanding of how figurative meanings are encoded in transformer representations but also provide a useful guidance for pinpointing specific processing failures in idiom comprehension.
+We present a multi-stage mechanism of idiom comprehension in causal transformers, by identifying distinct components and flows for figurative and literal interpretations.
+Causal interventions reveal (i) early retrieval of the idiomatic interpretation, (ii) immediate use of preceding context, with later refinement when it conflicts with the retrieved interpretation, (iii) a selective pathway that carries the figurative interpretation through intermediate layers, and a bypass route that favorably delivers the literal interpretation directly to the output.
+To sum up, these findings link representational dynamics to contextual disambiguation, yielding a concise mechanistic interpretation of idiom processing.
 Limitations
@@
 Moreover, knocking out or replacing the activations can lead the model to out-of-distribution behaviors and cast doubt on the robustness of any interpretability claim.
-Our study is limited to LLama3.2-1B-base model, which is much smaller than the current state-of-the-art transformers, scaling these analyses to larger models remains an open challenge.
-Also, we simplify idiom processing by excluding context, even though context critically shapes idiomatic meaning.
-Altogether, these factors limit the generalizability of our results.
+Moreover, we binarize idiom interpretation (figurative vs.
+literal), ignoring polysemy among figurative senses (e.g., ‘break the ice’ can be interpreted into initiate talk, ease tension, etc.), which can flatten nuance and bias evaluation.
+As a future direction, we can use soft labels for multiple figurative sense and distributional metrics.
 Ethics statement
@@
 Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102.
-Appendix A Selected top-20 attention heads
-We further extracted semantic heads $\mathcal{H}_{\mathrm{sem}}$, i.e., heads that show a positive $\Delta F(s_{a})$ and a negative $\Delta L(s_{a})$, which are processing ambiguous sentence in a literal way.
-Ablating them suppresses literal processing and restores figurative understanding.
-For comparison, we also randomly selected heads that are neither in $\mathcal{H}_{\mathrm{idiom}}$ nor $\mathcal{H}_{\mathrm{sem}}$.
-Type top-$20$ attention heads (layer, head)
-Idiomatic (0, 4), (0, 28), (1, 5), (2, 9), (0, 30), (0, 8), (0, 19), (2, 2), (1, 25), (1, 18), (4, 31), (1, 31), (0, 12), (0, 3), (2, 28), (2, 17), (2, 8), (2, 30), (1, 9), (4, 24)
-Semantic (1, 3), (3, 28), (3, 30), (0, 10), (7, 16), (1, 20), (2, 4), (3, 12), (3, 14), (6, 27), (2, 19), (10, 11), (4, 26), (0, 18), (6, 31), (9, 22), (2, 5), (0, 9), (1, 19), (0, 11)
+Appendix A Identified top-20 idiomatic heads
+Table 1 lists the attention heads most influential in retrieving figurative meanings while suppressing literal interpretations and randomly selected heads.
+Type top-2020 attention heads (layer, head)
+Idiomatic (0, 4), (1, 5), (0, 30), (0, 19), (0, 8), (9, 30), (2, 2), (1, 18), (0, 21), (2, 8), (1, 9), (4, 2), (1, 24), (1, 13), (3, 18), (0, 0), (3, 24), (0, 26), (1, 27), (0, 28)
 Random (9, 20), (10, 30), (2, 11), (2, 23), (1, 1), (9, 23), (13, 24), (8, 31), (3, 8), (13, 9), (10, 1), (5, 15), (7, 14), (7, 24), (4, 5), (14, 2), (9, 17), (4, 0), (14, 19), (6, 9)
-Table 1: Top-$20$ attention head sets for each type.
-Appendix B Activation divergence analysis between idiomatic and literal contexts
-To examine how idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ process idioma differently compared to literal contexts, we focus on their activations over the idiom token span.
-Let =$\mathcal{I}=\{i_{1},\dots,i_{K}\}\subseteq\{1,\dots,T\}$ be the indices of the idiom span tokens.
-Note that the literal paraphrase $s_{l}$ preserves the same idiom token span, hence $|\mathcal{I}|$ is identical for both $s_{a}$ and $s_{l}$.
-For each head $h\in\mathcal{H}_{\mathrm{idiom}}$, we compute the averaged activation from idiom span value vectors:
-$v_{a}\;=\;\frac{1}{\mathcal{I}}\sum_{i\in\mathcal{I}}\alpha^{(a)}_{h,i}\,V_{i},\;v_{l}\;=\;\frac{1}{\mathcal{I}}\sum_{i\in\mathcal{I}}\alpha^{(l)}_{h,i}\,V_{i},$
-where $\alpha^{(o)}_{h,i}$ and $\alpha^{(l)}_{h,i}$ are the attention weights on token $i$ in $s_{a}$ and $s_{l}$, respectively, and $V_{i}$ is the corresponding value projection.
-These mean vectors capture the direction in the head’s value subspace over the idiom span under each context.
-We then measure their alignment via cosine similarity:
-$\cos(v_{a},\,v_{l})=\frac{v_{a}\cdot v_{l}}{\|v_{a}\|\,\|v_{l}\|}.$
-A low cosine similarity between $v_{a}$ and $v_{l}$ means that the head’s representational direction shifts substantially when processing the idiomatic sentence $s_{a}$ compared to its literal paraphrase $s_{l}$, revealing its specialization in processing ambiguous sentence and extracting its figurative meaning.
-In contrast, a high cosine similarity indicates that the head maintains a similar representation across both contexts, reflecting a more general semantic role.
-For comparison, we also compute cosine similarities for the semantic heads $\mathcal{H}_{\mathrm{sem}}$ and for a random set of heads $\mathcal{H}_{\mathrm{rand}}$.
-For each sentence, we compute the cosine similarity $\cos(v_{a},v_{l})$ for every head in a given set of 20 attention heads, and then average these 20 values to obtain a single mean similarity (Figure 8).
-The cosine similarity across the dataset is shown in Figure 8.
-The $cos\_random$ ($M=0.87,\,SD=0.07$), $cos\_sem$ ($M=0.80,\,SD=0.09$) is significantly larger than $cos\_idiom$ ($M=0.76,\,SD=0.12$), ($p<0.05$).
-Figure 8: Cosine similarities between the attention activations on idiomatic sentences ($s_{a}$) and their literal paraphrase ($s_{l}$) for the top-20 idiomatic heads, semantic heads, and random heads.
-By confirming that the heads identified as specialized in retrieving the idioms’ figurative meaning indeed have divergent attention patterns when processing idioms compared to literal paraphrases, we strengthen the evidence that these heads are responsible for the unique processing required to interpret idiomatic expressions correctly.
-Appendix C Layer-wise interpretation flow through activation patching
-We additionally patch idiom’s because token activation into paraphrased sentences’ because token position. The conceptual details of the experiment is described as follow.
+Table 1: Top-2020 attention head sets.
+Appendix B Data generation pipeline
+We illustrate the data generation pipeline in Figure 9.
+Figure 9: The pipeline of data generation.
+Appendix C Layer-wise context disambiguation via logit lens
+If the MHSA layers 3-7 play a causal role in context disambiguation, then we would expect that the representation at the final idiom token is fully disambiguated towards either the literal or the figurative meaning, depending on the earlier context.
+We employ the logit lens method nostalgebraist (2020) to measure the log-probability assigned to figurative and literal interpretations in the figurative ($FC$) vs.
+literal context ($LC$) conditions.
+At each layer, the hidden representations are projected into the output vocabulary space, yielding candidate probabilities.
+We then measure the difference in log-probability mass between figurative and literal candidates across layers, separately isolating contributions from the MHSA and MLP components.
+(a) MLP FC-LC
+(b) MHSA FC-LC
+Figure 10: Layerwise differences in log-probability between figurative and literal candidates predictions under figurative (FC) vs. literal (LC) contexts, measured with the logit lens.
+ResultsFigure 10 shows a clear disambiguation effect for MHSA starting around layer 8.
+Once MHSA starts modulating activations in a context-dependent manner, it writes this information into the residual stream, which is subsequently processed by the MLP layers.
+In the early to middle layers (0–9) of the MLP, the difference in log probabilities between figurative and literal candidate predictions under figurative vs.
+literal contexts remains minimal.
+However, beginning at layer 10, MLP activations begin to diverge: idioms presented in figurative contexts are increasingly encoded with figurative interpretations, whereas idioms in literal contexts are increasingly encoded with literal interpretations.
+Appendix D Knockout for crucial role of early layers with other models
+Results across different models in Figure 11 consistently show that early layers retrieve idioms’ figurative interpretations while suppressing literal counterparts, unlike unambiguous sentences.
+Qwen2.5-0.5B
+(a) MLP
+(b) MHSA
+(c) MLP
+(d) MHSA
+Llama3.1-8B
+(e) MLP
+(f) MHSA
+(g) MLP
+(h) MHSA
+Qwen2.5-7B
+(i) MLP
+(j) MHSA
+(k) MLP
+(l) MHSA
+Figure 11: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for sentences $s\in\{s_{a},s_{f},s_{l}\}$.
+Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals.
+X-axis: Layers.
+Gray dashed line: $\Delta I=0$ (no effect).
+Red asterisk (*): Significant difference between $\Delta F(s_{a})$ and the others (paired $t$-test, $p<0.05$).
+The difference at * marked layer is larger than the average difference across all layers.
+Appendix E Activation patching for information flow with other models
+Across models, the intermediate and direct pathways consistently prefer different interpretations (see Figure 12).
+Qwen2.5-0.5B
 […]
 […]
-Figure 9: Description of the activation patching experiment for identifying interpretation flow.
-L represents literal interpretation; F represents figurative interpretation.
-(a) By replacing ‘because’ token $s_{f}$ activation with $s_{a}$, we expect literal interpretation to be added in the intermediate channel while keeping the direct route flow to be activated with naive figurative meaning.
-(b) By replacing ‘because’ token $s_{l}$ activation with $s_{a}$, we expect figurative interpretation to be added via intermediate path, while keeping both paths for literal interpretation.
 […]
 […]
-Figure 10: Layer-wise interpretation shift $\Delta I(s_{\mathrm{x}}\hookleftarrow s_{a}),\,\mathrm{x}\in\{f,l\}$ after activation patching on the ‘because’ token across three patch directions.
-Y-axis: Mean values of $\Delta L(s_{\mathrm{x}}\hookleftarrow s_{a})$, $\Delta F(s_{\mathrm{x}}\hookleftarrow s_{a})$.
+Llama3.1-8B
+[…]
+[…]
+[…]
+[…]
+Qwen2.5-7B
+[…]
+[…]
+[…]
+[…]
+Figure 12: Layer-wise interpretation shift after patching in activations from $s_{f}$ and $s_{l}$ and vice versa.
+The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
+Appendix F Knockout for context disambiguation with other models
+Across different models, Figure 13 consistently indicate that early MLP layers retrieve idioms’ figurative meanings, and following layers promotes contextual cues via MHSA when a literal context is present.
+Qwen2.5-0.5B
+(a) MLP FC
+(b) MHSA FC
+(c) MLP LC
+(d) MHSA LC
+Llama3.1-8B
+(e) MLP FC
+(f) MHSA FC
+(g) MLP LC
+(h) MHSA LC
+Qwen2.5-7B
+(i) MLP FC
+(j) MHSA FC
+(k) MLP LC
+(l) MHSA LC
+Figure 13: Sublayer-wise interpretation shift $\Delta I(C+s_{a})$ after ablating activations at idiom span, for contexts $C\in\{FC,LC\}$.
+Y-axis: Mean values of $\Delta L(C+s_{a})$, $\Delta F(C+s_{a})$ with 95% confidence intervals.
 X-axis: Layers.
 Gray dashed line: $\Delta I=0$ (no effect).
-Red asterisk (*): Significant shift between literal ($C_{l}$) and figurative ($C_{f}$) candidate probability (paired $t$-test, $p<0.05$).
-As expected, by adding idiom’s interpretation via intermediate channel, since it’s adding literal interpretation, there’s a increase in $\Delta L(x_{f}\hookleftarrow s_{a})$ by $0.1$, whereas $\Delta F(x_{f}\hookleftarrow s_{a})$ drops by $-0.07$ across layers (Figure 10(a)).
-Whereas, when adding the idiom’s figurative interpretation into intermediate path, there’s a big increase in $\Delta F(s_{l}\hookleftarrow s_{a})$ by $0.26$ at layer 7 and decrease in $\Delta L(s_{l}\hookleftarrow s_{a})$ by $-0.17$.
-This implies that when there’s idiomatic representation passing the intermediate path, to make final prediction as close to figurative one, the model tries to amplify figurative meaning that is passing the intermediate path so that the total amount of activations for figurative interpretation to be same, which should be more than its literal one.
-That is, injecting the idiom’s literal activation into the intermediate channel increases $\Delta L(x_{f}\hookleftarrow s_{a})$ by up to $0.1$ in the mid-layers (4-9).
-Conversely, injecting figurative activation raises $\Delta F(s_{l}\hookleftarrow s_{a})$ by $0.26$.
-This asymmetric increase indicates that when an idiomatic representation passes the intermediate path, the model actively amplifies figurative interpretation so it dominates the final prediction.
-Appendix D Analyzing edge-level intermediate channel through attention knockout
-To trace the flow of information, we perform fine-grained intervention on the MHSA sublayers, since they mediate communication between token positions, therefore, any critical information must pass through them Geva et al.
-(2023).
-To locate the intermediate and direct path for each interpretation, we focus on the attention scores between idiom to subsequent token (i.e., because) for the intermediate path, and the idiom to last token of the sentence for the direct path findings.
-We investigate how this interpretation of flow is done internally by knockout the attention edges between target tokens.
-When critical attention edges for a certain interpretation of the idiom are blocked, it will result in severe degradation in probabilities of certain candidates.
-Therefore, we test whether critical information propagates between two hidden representations at a specific layer, by zeroing-out all the attention edges between them.
-Formally, let $i,j\in\{1,\dots,T\}$ with $j\leq i$ be two positions in the sequence.
-To block token $i$ (query) from attending to token $j$ (key) at layer $\ell<L$, we overwrite the corresponding entries in the causal self-attention mask $M$ at layer $\ell+1$ across all heads $h\in\{1,\dots,H\}$:
-$M_{i,\,j}^{\ell+1,\,h}\;=\;-\,\infty\quad\forall\,h\in\{1,\dots,H\}.$
-(a) Idiom $\rightarrow$ Subsequent
-(b) Idiom $\rightarrow$ Last
-Figure 11: Layer-wise interpretation shift $\Delta I(s_{a})$ after knock out the attention edges between tokens.
-Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$.
-X-axis: Layers.
-Gray dashed line: $\Delta I=0$ (no effect).
-Red asterisk (*): Significant shift between literal ($C_{l}$) and figurative ($C_{f}$) candidate cumulative probabilities (paired $t$-test, $p<0.05$).
-This restricts the query position from obtaining information from the key position, at that particular layer ($\ell+1$).
-For each layer, we block attention edges between (a) idiom span tokens $\rightarrow$ subsequent token (b) idiom span tokens $\rightarrow$ last token.
-If the specific path is important for a specific interpretation, then there should be significant drop of $I(s_{a})$ value.
-Figure 11 shows the result.
-In Figure 11(a) at layer 5, it shows that there’s a drop in the $\Delta F(s_{a})$ by $-0.02$, while increase in $\Delta L(s_{a})$, by $0.02$.
-This shows that ablating attention from the idiom span to ‘because’ token disrupts the indirect path that would do disambiguation process where it suppresses the literal one.
-Without the idiom attending to the ‘because’ token, the model can no longer effectively disambiguate figurative meaning from literal meaning.
-At layer 11 and 13, as in Figure 11(b), $\Delta L(s_{a})$ declines while $\Delta F(s_{a})$ rises.
-This divergence indicates that ablating the direct idiom to last token pathway prevents the model from passing the literal interpretation.
+Red asterisk (*): Significant difference between $\Delta F(C+s_{a})$ and the others (paired $t$-test, $p<0.05$).
+The difference at * marked layer is larger than the average difference across all layers.
 Ambiguous sentence Literal paraphrase Figurative paraphrase Literal candidates Figurative candidates
@@
 They let him off the hook because he was a They removed him from fishing tackle because he was a They freed him from responsibility because he was a threat, bad, convicted, little, bit, danger, sexual, fish, ped, white, criminal, racist, bully, black, terrorist, thief, political, jerk, rap, big minor, child, foreign, slave, kid, good, mad, juvenile, victim, boy, stranger, young, youth, mere, student, fool, first, teenager, prisoner, friend
-She was looking for a needle in a haystack because she was a She was searching for a neddle within a straw because she was a She was facing an impossible search because she was a needle, craft, seam, straw, tiny, hay, detective, master, haystack, witch, farmer, pro, camel, professional, busy, crazy, giant, chicken, cow, neat woman, victim, single, young, black, female, girl, slave, mother, stranger, new, first, non, private, white, novice, minority, prisoner, foreign, mom
+She was looking for a needle in a haystack because she was a She was searching for a needle within a straw because she was a She was facing an impossible search because she was a needle, craft, seam, straw, tiny, hay, detective, master, haystack, witch, farmer, pro, camel, professional, busy, crazy, giant, chicken, cow, neat woman, victim, single, young, black, female, girl, slave, mother, stranger, new, first, non, private, white, novice, minority, prisoner, foreign, mom
 Table 3: Biased data examples reflecting stereotypes embedded in the language model.

### v5

@@
 (iii) Then, selective, competing pathways carry both interpretations: an intermediate pathway prioritizes the figurative interpretation and a parallel direct route favors the literal interpretation, ensuring that both readings remain available.
-Our findings provide mechanistic evidence for idiom comprehension in autoregressive transformers [^1] Code and dataset are available at https://github.com/sori424/idiom_processing under the MIT License..
+Our findings provide mechanistic evidence for idiom comprehension in autoregressive transformers [^1] Code and dataset are available at https://github.com/sori424/idiom_processing.
 […]
@@
 The tokens in each set are mutually exclusive and chosen for their high likelihood of association with the respective interpretation.
-Below, we outline the dataset construction process, and see Figure 9:
+Below, we outline the dataset construction process, and see Figure 9.
 1.
@@
 (2024).
-The sentences below are three examples of the generated ambiguous sentence ($s_{a}$), figurative paraphrase ($s_{f}$), and literal paraphrase ($s_{l}$).
+The sentences below are three examples of the generated ambiguous sentence ($s_{a}$), figurative paraphrase ($s_{f}$), and literal paraphrase ($s_{l}$):
 […]
@@
 3.1 Information flow in transformers
-Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{T}]$ over a vocabulary $V$, each token is embedded as $\mathbf{x}_{i}^{0}=\mathrm{PE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{PE}$ denotes a positional encoding function that injects position $i$ into the token embedding Vaswani et al.
+Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{N}]$ over a vocabulary $V$, each token is embedded as $\mathbf{x}_{i}^{0}=\mathrm{PE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{PE}$ denotes a positional encoding function that injects position $i$ into the token embedding Vaswani et al.
 (2017), and $d$ is the model’s hidden dimension.
@@
 4.1 Probing sublayers through knockout
-To locate the mechanisms for retrieving figurative interpretations from the model’s parameters, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$.
+To locate the mechanisms for retrieving figurative interpretations, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$.
 By knocking out the activations across each layer, we assess how the probabilities of predicting specific candidate tokens ($\Delta F(s_{a}),\,\Delta L(s_{a})$) change.
-To narrow down the points to which we can ascribe a functional role for a specific computation component, we limit the intervention to the MLP and MHSA in each block.
+To narrow down the points to which we can ascribe a functional role for a specific computation component, we limit the intervention to a MLP and MHSA in each block.
 (a) MLP
@@
 Figure 2: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for sentences $s\in\{s_{a},s_{f},s_{l}\}$.
-Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals.
+Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals across idiom sentences.
 X-axis: Layers.
 Gray dashed line: $\Delta I=0$ (no effect).
-Red asterisk (*): Significant difference between $\Delta F(s_{a})$ and the others (paired $t$-test, $p<0.05$).
-The difference at * marked layer is larger than the average difference across all layers.
-Results Figure 2(a) shows by how much the model’s interpretation of the idiom changes when an MLP sublayer is knocked out.
+The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
+Results Figure 2(a) shows by how much the model’s interpretation of the idiom changes when the MLP sublayer is knocked out.
 We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2).
@@
 This indicates that early layer MLPs are causally important for reading out the figurative interpretation while potentially suppressing the literal one.
-For later layers ($\geq$ 4), MLP sublayer knockout does not affect the model’s interpretation, where the corresponding patching effect converges to zero.
+For later layers ($\geq$ 4), the knockout does not affect the model’s interpretation, where the corresponding patching effect converges to zero.
 Figure 2(b) shows the effect of MHSA knockout, which is similar to what we observed with MLP sublayer knockout.
@@
 We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$.
+To operationalize this, we define idiomatic heads ($\mathcal{H}_{\mathrm{idiom}}$) as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$.
 (a) $\Delta F(s_{a})$
@@
 Figure 3: Heatmaps of the (a) $\Delta F(s_{a})$ (b) $\Delta L(s_{a})$ when ablating individual attention heads at the idiom span.
-Idiomatic heads: Heads those are crucial for retrieving the figurative interpretation of idiom; $\Delta F(s_{a})$ and $\Delta L(s_{a})$.
-To operationalize this, we define idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$.
-Figure 3 shows that there is a subset of attention heads that simultaneously exhibits large $\Delta F(s_{a})$ (Figure 3(a)) and $\Delta L(s_{a})$ (Figure 3(b)).
+Idiomatic heads: Heads those are crucial for retrieving the figurative interpretation of idiom; $\Delta F(s_{a})$ and $\Delta L(s_{a})$, simultaneously.
+We identify attention heads whose removal reliably decreases figurative probability ($\Delta F(s_{a})$), while increasing literal probability ($\Delta L(s_{a})$) for the idiom sentences.
 To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples for each head’s contribution to the interpretation.
 We then rank all heads by, on the one hand, $\Delta F(s_{a})$, and, on the other hand, $\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders (selected attention heads listed in Appendix Table 1).
+Figure 3 shows that there is a subset of attention heads that simultaneously exhibits large $\Delta F(s_{a})$ (Figure 3(a)) and $\Delta L(s_{a})$ (Figure 3(b)).
 4.3 Causal role of MLP sublayers and idiomatic heads in figurative reading
-To validate that the early MLP sublayers and the selected idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ are critical for the figurative reading, we patch the activations on $s_{a}$ (at idiom span) into the corresponding components when processing $s_{l}$.
-As a control, we patch randomly chosen MLP sublayers and heads.
+To show that the components that we found (i.e., early MLP sublayers and $\mathcal{H}_{\mathrm{idiom}}$) are causally sufficient for figurative reading, we patch the activations on $s_{a}$ at idiom span tokens into the corresponding components when processing $s_{l}$: patching their activations from the idiom sentence into the literal paraphrase increases $\Delta F(s_{l}\hookleftarrow s_{a})$ and decreases $\Delta L(s_{l}\hookleftarrow s_{a})$, while patching random components has no effect.
 ResultsPatching these components boosts the figurative interpretation, as quantified by $\Delta F(s_{l}\hookleftarrow s_{a})$ (M=0.17,SD=0.49M=0.17,\,SD=0.49) while suppressing the literal interpretation, $\Delta L(s_{l}\hookleftarrow s_{a})$ (M=0.29,SD=0.58M=-0.29,\,SD=0.58).
@@
 […]
+[…]
+[…]
 Figure 5: Conceptual description of the activation patching experiments for tracing information flow (L = literal interpretation; F = figurative interpretation).
+[…]
+[…]
 […]
@@
 This indicates that the figurative interpretation is suppressed in favor of the literal interpretation.
-We interpret this as showing that a figurative injection doesn’t lead to suppression of the literal interpretation (an asymmetry between increases and decreases).
+The asymmetry between increases and decreases show that a figurative injection doesn’t lead to suppression of the literal interpretation.
 By contrast, a literal injection must first suppress the figurative interpretation in that subspace to have an effect, producing a more symmetric push–pull tradeoff.
@@
 These results indicate that the direct path predominantly conveys literal information.
-The pattern is consistent across different models (Appendix E).
+This pattern is consistent across different models (Appendix E).
 6 Localizing idiom disambiguation in context
@@
 (d) MHSA LC
-Figure 7: Sublayer-wise interpretation shift $\Delta I$ after knockout activations at idiom span, for contexts FCFC: Figurative Context; LCLC: Literal Context.
+Figure 7: Sublayer-wise interpretation shift $\Delta I$ after knockout activations at idiom span with contexts.
+The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
 ResultsFigure 7(a) shows that for idioms in a figurative context, knocking out the MLP in the early layers (0–3) results in a significant drop in $\Delta F$ (min: 0.62-0.62) and increase in $\Delta L$ (max: $0.66$), consistent with the pattern observed for idioms without context (Figure 2(a)).
@@
 This result is consistent across different models (Appendix F).
-For idioms in a literal context (7(b)), ablating the MLP layer 1 component leads to a significant drop in $\Delta F$ (min: 0.33-0.33) and an increase in $\Delta L$ (max: $0.09$), indicating that the figurative interpretation is still being retrieved.
+For idioms in a literal context (7(b)), ablating the MLP layer 1 leads to a significant drop in $\Delta F$ (min: 0.33-0.33) and an increase in $\Delta L$ (max: $0.09$), indicating that the figurative interpretation is still being retrieved.
 The effect persists but is small compared with the figurative context condition, where early layers at the idiom token appear to encode context that suppresses figurative retrieval while biasing toward the literal interpretation.
@@
 This indicates that, after the initial retrieval of figurative interpretation, the literal interpretation, which is disambiguated from context, is refined through mainly MHSA mediated integration with surrounding tokens.
-In sum, the results point to a two stage context disambiguation: (a) early layers retrieve figurative evidence while beginning to encode context, followed by (b) mid-layers, MHSA-driven contextual disambiguation that resolves conflicts in favor of the context-appropriate interpretation.
-6.2 Probing MHSA-mediated contextual disambiguation through KVQ patching
-To examine in detail how MHSA mediates contextual disambiguation in early-mid layers, we perform KVQ patching, in which MHSA’s key–value (KV) activations are replaced while selectively preserving idiom tokens’ query (Q) activations across layers, thereby steering the model’s attention toward alternative contexts.
-That is, we redirect idiom token span’s attention from a figurative-context to the literal context, and vice versa.
-If the model’s final prediction changes from figurative to literal (or vice versa) after patching, it demonstrates that the MHSA mechanism in those specific layers is responsible for using the context to steer the disambiguation process.
-[…]
-[…]
-[…]
+In sum, the results point to a two stage context disambiguation: (1) early layers retrieve figurative evidence while beginning to encode context, followed by (2) mid-layers, MHSA-driven contextual disambiguation that resolves conflicts in favor of the context-aligned interpretation.
+6.2 Probing MHSA contextual disambiguation with Query-preserved KV patching
+To examine in detail how MHSA mediates contextual disambiguation in early-mid layers, we perform Query-preserved Key-Value patching, in which MHSA’s Key–Value (KV) activations are swapped while selectively preserving idiom tokens’ query (Q) activations across layers, thereby redirecting the model’s attention toward alternative contexts.
+That is, we switch idiom token span’s attention from a figurative-context to the literal context, and vice versa.
+If the model’s final prediction changes from figurative to literal (or vice versa) after patching, it demonstrates that the MHSA mechanism in those specific layers is responsible for using the context in the disambiguation process.
+(a) FC $s_{a} \hookleftarrow$ LC $s_{a}$
+at idiom span
+(b) LC $s_{a} \hookleftarrow$ FC $s_{a}$
+at idiom span
+Figure 8: Layer-wise interpretation shifts $\Delta I$ under Q-preserved KV patching.
+Y-axis: Mean $\Delta L(s_{a})$ and $\Delta F(s_{a})$ across idiom sentences.
+X-axis: Layers.
+Gray dashed line: $\Delta I=0$ (no effect).
+The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
 ResultsAs shown in Figure 8, patching in KVs with alternative context yields a shift in interpretation of idiom.
@@
 Causal interventions reveal (i) early retrieval of the idiomatic interpretation, (ii) immediate use of preceding context, with later refinement when it conflicts with the retrieved interpretation, (iii) a selective pathway that carries the figurative interpretation through intermediate layers, and a bypass route that favorably delivers the literal interpretation directly to the output.
+Considering idioms as multiword expressions, our results suggest that early layers (especially MLPs) perform a “detokenization” step Gurnee et al.
+(2023); Elhage et al.
+(2021) that compresses the idiom into a unified figurative representation, while later pathways can “retokenize” it into a competing literal interpretation.
 To sum up, these findings link representational dynamics to contextual disambiguation, yielding a concise mechanistic interpretation of idiom processing.
@@
 As a future direction, we can use soft labels for multiple figurative sense and distributional metrics.
-Ethics statement
+Ethical Considerations
 One of the intended uses of the Llama-3.3-70B-instruct model is content generation, which aligns with our use of the model for generating the candidates for next token prediction task Grattafiori et al.
@@
 Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102.
+We gratefully acknowledge the stimulating research environment of the GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under project number 471607914.
 Appendix A Identified top-20 idiomatic heads